### v1

# Interaction-Aware Model Predictive Decision-Making for Socially-Compliant Autonomous Driving in Mixed Urban Traffic Scenarios

Abstract
This paper presents the experimental validation of an interaction-aware model predictive decision-making (IAMPDM) approach in the course of a simulator study. The proposed IAMPDM uses a model of the pedestrian, which simultaneously predicts their future trajectories and characterizes the interaction between the pedestrian and the automated vehicle. The main benefit of the proposed concept and the experiment is that the interaction between the pedestrian and the socially compliant autonomous vehicle leads to smoother traffic. Furthermore, the experiment features a novel human-in-the-decision-loop aspect, meaning that the test subjects have no expected behavior or defined sequence of their actions, better imitating real traffic scenarios. Results show that intention-aware decision-making algorithms are more effective in realistic conditions and contribute to smoother traffic flow than state-of-the-art solutions. Furthermore, the findings emphasize the crucial impact of intention-aware decision-making on autonomous vehicle performance in urban areas and the need for further research.
keywords:
Human-Machine Interaction , Human in the Loop , Pedestrian Motion Prediction , Simulator Experiment , Autonomous Vehicles , Model Predictive Controller
journal: Elsevier
[…]
[label0] addressline=balint.varga2@kit.edu, organization=Institute of Control Systems, Karlsruhe Institute for Technology,city=Karlsruhe, postcode=D-76131, country=Germany
[…]
[label1] organization=Wuerzburg Institute for Traffic Sciences GmbH, city=Veitshoechheim, postcode=D-97209, country=Germany
[…]
[label2] organization=Faculty of Engineering - Mechanical Engineering Dept, University of Alberta, city=Edmonton AB, postcode=Ca-T6G 2H5, country=Canada
1 Introduction
Highly automated and autonomous vehicles (AVs) are increasingly becoming a part of our daily lives (Elliott et al., 2019)(Zhou et al., 2022). Integrating these systems into society largely depends on the trust of vulnerable road users, such as cyclists and pedestrians, whose safety is essential (Bouzidi and Hashemi, 2023). Incidents involving automated driving functions often make headlines, fostering public skepticism towards these technologies. Consequently, significant research efforts are directed at enhancing automated vehicles with advanced communication channels and decision-making algorithms designed to manage complex scenarios. For example, in urban environments and city centers – where vehicles move at slower speeds, and pedestrians may cross unexpectedly – an effective human-machine interaction is critical to fostering trust in these systems.
Figure 1: The pedestrian crosses the street at an unsignalised intersection interacting with an autonomous vehicle. With the courtesy of version1 GmbH.
In recent years, advanced algorithms have enabled automated vehicles to interact with pedestrians and cyclists (Hu et al., 2018), (Razmi Rad et al., 2020). The efficiency of these algorithms from the literature is often analyzed either in pure simulation setups (Kampitakis et al., 2023) or in human-in-the-loop tests see e. g. Brand et al. (2024). These human-in-the-loop experiments [^1] Note that the terms experiment and study are used interchangeably in the course of this paper. are usually conducted in virtual environments (Razmi Rad et al., 2020) (Luu et al., 2022). However, literature-based experiments restrict decision possibilities: the vehicle’s behaviors or the test subjects’ expected reactions are predefined, making the experiment more controlled and analysis more accessible, limiting their application in real-world scenarios.
Therefore, this work validates an interaction-aware model predictive decision-making (IAMPDM) within a simulator experiment using a human-in-the-decision-loop setup, where test subjects’ reactions are unspecified to replicate real-world scenarios better. Since field studies with actual vehicles can be risky, human-in-the-loop simulations offer a safer alternative, see e.g Razmi Rad et al. (2020), Tran et al. (2021), Bindschädel et al. (2021) or Kalantari et al. (2023). This paper provides a human-in-the-decision-loop simulator study for the first time with the following contributions:
  • 1.
    Adapting the IAMPDM for the real-time experiments,
  • 2.
    Designing a suitable experimental human-in-the-decision-loop simulator study, and
  • 3.
    Conducting the experiment with 25 participants, offering analysis and recommendations for the real-world implementation of the proposed IAMPDM.
The paper is structured as follows: In Section 2, the state of the art of intention-aware automated vehicles is presented. Section 3 presents the concept of the IAMPDM. The simulator setup of the validation and the experimental design are given in Section 4. Section 5 provides the results of the experiment and a further discussion about the limitations. Finally, the paper is summarized in Section 6.
2 State of the Art
This section provides a short comprehensive literature review of interaction-aware decision-making algorithms, which became large in recent years, since the topic is getting more and more relevant. Therefore, the following state-of-the-art overview focuses on
  • 1.
    decision-making methods, models and algorithms [^2] For simplicity, the terms decision-making methods, models, and algorithms will collectively be referred to as decision-making concepts. and
  • 2.
    experiments and studies, which can verify and validate these concepts.
For more detailed reviews, the reader is referred to Predhumeau et al. (2023).
2.1 Decision-Making Concepts
In the literature, decision-making concepts can be categorized into two main groups: model-based and model-free approaches.
2.1.1 Model-free Approaches
Model-free concepts do not utilize physics-based or logic-based models. Instead, they learn optimal behaviors through trial and error by directly interacting with the environment. These methods use environmental feedback to refine decision-making, often employing algorithms such as Q-learning or policy gradient methods. Generally, model-free approaches are well-suited for complex scenarios, such as modeling human behavior. A review on this topic can be found in Di and Shi (2021).
In Russo et al. (2021), a Reinforcement Learning (RL) approach is presented addressing pedestrian collision avoidance in autonomous driving systems. The focus is on managing unexpected pedestrian crossings and tracking a specific trajectory. The Deep Deterministic Policy Gradient (DDPG) algorithm is utilized to learn continuous actions. The paper discusses the reward function, agent architecture, and environment model used for training and testing the DDPG-based agent through numerical simulations. Similar, data-based concepts are presented in Raghuram Kadali et al. (2014), Rezaee et al. (2021), Trumpp et al. (2022) and Zhang et al. (2022).
However, the main drawbacks of these model-free concepts are that they cannot easily ensure safety or guarantee the reliability needed in high-risk environments, such as AV-pedestrian interactions. This limitation means that pure learning-based systems are not mature enough to be safely used in real traffic situations. It is to be expected that model-free decision-making approaches will become easier to validate and verify for traffic admission, for which one typical method is the combination of model-free and model-based elements as discussed in Korbmacher and Tordeux (2022) or Derajić et al. (2024).
2.1.2 Model-based Approaches
On the other hand, model-based concepts utilize mathematical models to describe AV-pedestrian interactions using logical or physical principles. These methods rely on this predefined model to predict future states and make decisions. Furthermore, they can handle constraints and uncertainties in the system-design phase, leading to improved safety and reliability. More detailed overviews of model-based methods can be found in Camara et al. (2021), Gulzar et al. (2021) or Predhumeau et al. (2023).
In Domeyer et al. (2022), a perceptual model is established based on the geometrical relations between drivers and pedestrians, emphasizing the implications for vehicle automation. It explores the coupling of driver and pedestrian. However, only a simulation analysis is provided, with no strong indications for real-world usage. Game theoretical models are presented in Cheng et al. (2015), Skugor et al. (2020) and Pavelko et al. (2022). In these works, the interaction between the pedestrian and vehicle is modeled as a game with two players who optimize their own objective functions by their joint actions. These models can model the interaction between the two players, but they are not suitable to predict their future joint actions.
On the other hand, in Chen et al. (2023), a model predictive control (MPC) formulation is proposed, which can predict the future behavior of the pedestrian. In Škugor et al. (2023), the model predictive formulation is extended with a stochastic component, providing a more realistic overall behavior of the proposed algorithm.
However, these concepts often have a high computational complexity, therefore they are often not suitable for real-time application on a state-of-the-art automotive hardware.
2.2 Validation of Decision-Making Concepts
Since the main contribution of the work is the experimental study with a human-in-the-decision-loop character, the aftercoming part of the overview focuses on pedestrian-autonomous vehicle experiments from the literature. Reviews on the virtual reality studies of pedestrian-automated vehicle interaction can be found in Tran et al. (2021) Zhou et al. (2022).
The two main focuses in the literature are the impact of the external human-machine interface (eHMI) on trust and safety. For instance, in Prattico et al. (2021), various eHMI concepts are compared, and the influence of the eHMI on the test subjects’ trust is investigated. Similarly, in Feng et al. (2023), the impact of eHMI on pedestrian behavior at road crossings is analyzed in a virtual reality setting. The variation of the communication strategies of the automated vehicle is analyzed in Löcken et al. (2019). The authors proposed five groups for the different communication categories, which are implemented and tested in a virtual reality setup. The results showed that the higher the complexity, the more the test subjects trusted the technical system.
The study in Yang et al. (2024) focuses on how pedestrians’ head movements should be interpreted in the presence of automated vehicles at a virtual crossroad, offering insights into non-verbal cues and communication patterns in pedestrian-vehicle interaction for enhanced safety. However, the test subjects in this study are not in the decision loop, which means that the interaction between humans and automation has a negligible effect on the study results.
Thus, the works from the state of the art provide a wide range of experiments and investigations, however, none of these focus on the effect of various decision-making algorithms on the preference of the pedestrians.
2.3 Shortcoming of the State-of-the-Art Methods
The decision-making and control algorithms from the state of the art have shortcomings that 1) they are either too complex for real-time implementation on general automotive hardware or 2) they are hard to validate in order to pass motor vehicle registration of AVs.
On the other hand, experimental studies from the literature usually use predefined strategies for autonomous vehicles in the case of human-in-the-loop investigations. However, human-in-the-decision-making-loop is an inevitable extension for the further analysis of human factors. This means that the experiments should have no predefined behavior, which can have an impact on the results of the decision-making.
Therefore, our work addresses these challenges and provides an IAMPDM that can run in real time. Furthermore, this work also provides an investigation having a human-in-the-decision-making-loop character.
3 Adjustment of the Model Predictive Algorithm
3.1 Model Description
The IAMPDM implements an explicit motion model for the pedestrians’ motion. This is necessary to make the decision-making algorithm of the AV more suitable for city-center traffic scenarios with low driving speeds. The core idea of this IAMPDM is presented in Varga et al. (2023b) that we used for our implementation; and we validated the model using the data from Yang et al. (2019).
The proposed model uses a sigmoid function based on the time-to-collision (TTC) values to model the probability of the pedestrian’s velocity changes. In the following, a detailed and more computationally efficient formulation is given. For the explicit motion model, it is assumed that the pedestrian’s choice of speed at the next time step can be modeled by
$\dot{y}_{\mathrm{ped}}(t)=\frac{1}{1+\exp\left(-TTC(t)+c\right)}\cdot v_{\mathrm{ped}}^{\mathrm{ref}},$ (1)
where c is a parameter that can be used to adjust the pedestrian’s characteristics, such as cautious or aggressive behavior. The $TTC$ is the time to collision and is computed by
$TTC(t)=\frac{x_{\mathrm{ped}}(t)-x_{\mathrm{veh}}(t)}{v_{\mathrm{veh}}(t)}-\frac{y_{\mathrm{veh}}(t)-y_{\mathrm{ped}}(t)}{v_{\mathrm{ped}}^{\mathrm{ref}}}.$ (2)
The output function of (1) is a general sigmoid function and ranges between 0 and 1, which can be treated as the probability of the pedestrian crossing. The greater the $TTC$ value, the more likely it is that the pedestrian would choose to cross at a reference speed. Therefore, only a reference speed of the pedestrian $v_{\mathrm{ped}}^{\mathrm{ref}}$ needs to be identified for the model. This model allows for real-time prediction of pedestrian motion and model-based interaction between the pedestrian and automated vehicle, making it simultaneously suitable for practical applications.
The validity of (1) can be supported by the research of Kadali and Vedagiri (2020) and Forde and Daniel (2021), which propose also velocity-based models for pedestrian-vehicle interactions. The main benefit of the proposed model is that the decision layer is integrated into to control layer, thus the MPC can solve the trajectory planning and decision making tasks in one step.
Gap distance model (see e.g. (Tian et al., 2024)) inherently assumes that the pedestrian’s decision-making is based on the current gap value and assessing the safety without considering how that gap might change in the future. On the other hand, (1) modulates the interaction dynamics between pedestrian and vehicle, which evolves over time. It can account for variables such as speed, acceleration, and changing distances in a continuous manner and used for prediction interaction dynamics.
Assuming a linear dynamics of the vehicle [^3] This assumption is commonly utilized by the application of autonomous vehicles, see e.g. Lynch and Park (2017, Chapter 13). and that the pedestrian moves in the $y$ direction, the following discrete dynamic system is obtained:
$\underbrace{\begin{bmatrix}x_{\mathrm{veh}}(t+\Delta t)\\ \dot{x}_{\mathrm{veh}}(t+\Delta t)\\ y_{\mathrm{ped}}(t+\Delta t)\\ \dot{y}_{\mathrm{ped}}(t+\Delta t)\end{bmatrix}}_{\boldsymbol{x}(t+\Delta t)}=\underbrace{\begin{bmatrix}1&{\Delta}t&0&0\\ 0&1&0&0\\ 0&0&1&\Delta t\\ 0&0&0&0\end{bmatrix}}_{{\mathbf{A}}}\underbrace{\begin{bmatrix}x_{\mathrm{veh}}(t)\\ \dot{x}_{\mathrm{veh}}(t)\\ y_{\mathrm{ped}}(t)\\ \dot{y}_{\mathrm{ped}}(t)\end{bmatrix}}_{\boldsymbol{x}(t)}$
$+\underbrace{\begin{bmatrix}0.5\cdot{\Delta}t^{2}\\ \Delta t\\ 0\\ 0\end{bmatrix}}_{{\mathbf{B}}}u_{\mathrm{veh}}(t)+\underbrace{\begin{bmatrix}0\\ 0\\ 0\\ \frac{v_{\mathrm{ped}}^{\mathrm{ref}}}{1+e^{-TTC(t)+c}}\end{bmatrix}}_{\boldsymbol{z}(t)}$ (3)
where the desired acceleration of the vehicle ${u_{\mathrm{veh}}(t)=a_{\mathrm{des}}}$ is the system input. With the model (3.1), the prediction of the future states for $N$ steps is possible using the batch formulation Borrelli et al. (2017):
$\mathcal{A}=\begin{bmatrix}\mathbf{A}\\ \mathbf{A}^{2}\\ \vdots\\ \mathbf{A}^{N}\end{bmatrix}\hskip 8.53581pt\mathrm{and}\hskip 8.53581pt\mathcal{B}=\begin{bmatrix}\mathbf{B}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{AB}&\mathbf{B}&\ldots&\mathbf{0}\\ \vdots&\ddots&\ddots&\vdots\\ \mathbf{A}^{N-1}\mathbf{B}&\ldots&\mathbf{AB}&\mathbf{B}\end{bmatrix}\hskip 5.69054pt\mathrm{and}\hskip 8.53581pt\mathcal{Z}=\begin{bmatrix}\mathbf{1}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{A}&\mathbf{1}&\ldots&\mathbf{0}\\ \vdots&\ddots&\ddots&\vdots\\ \mathbf{A}^{N-1}&\ldots&\mathbf{A}&\mathbf{1}\end{bmatrix}.$
The prediction of the future state and input vectors are
${\mathbf{x}}_{s}=\mathcal{A}{\mathbf{x}}_{0}+\mathcal{B}{\mathbf{u}}_{s}+\mathcal{Z}{\mathbf{z}}_{s},$ (4)
where for the sake of simplicity the index $s$ is used for the vector sequences: ${\mathbf{x}}_{s}=[{\mathbf{x}}(t),\>{\mathbf{x}}(t+\Delta t)\>\ldots\>{\mathbf{x}}(t+\Delta\cdot(N-1)t)]^{T}$, for the future system states, $\boldsymbol{u}_{s}=[u(t),\>u(t+\Delta t)\>\ldots\>u(t+\Delta\cdot(N-1)t)]^{T}$ for the optimizing future inputs and ${\mathbf{z}}_{s}=[{\mathbf{z}}(t),\>{\mathbf{z}}(t+\Delta t)\>\ldots\>{\mathbf{z}}(t+\Delta\cdot(N-1)t)]^{T}$ for the pedestrian future dynamics.
3.2 Cost Function Formulation
In order to formulate an MPC, the cost function with three components is defined
$J_{\mathrm{MPC}}=J_{\mathrm{com}}+J_{\mathrm{ref}}+J_{\mathrm{safe}},$ (5)
where three objectives are formulated: 1) comfort of the vehicle, 2) holding the reference velocity of the vehicle, and 3) safety of interaction between pedestrian and AV. These quadratic cost functions are
Figure 2: Bird’s eye view of the scenario with the relevant distances for decision-making
$J_{\mathrm{com}}$ $=\mathrm{w}_{\mathrm{com}}\cdot\boldsymbol{u}^{2}_{s}$ (6a)
$J_{\mathrm{ref}}$ $=\boldsymbol{x}^{\mathsf{T}}_{s}(t)\cdot\mathcal{Q}_{\mathrm{ref}}\cdot\boldsymbol{x}_{s}(t)$ (6b)
$J_{\mathrm{safe}}$ $=\mathrm{w}_{\mathrm{safe}}\cdot\frac{1}{\boldsymbol{x}^{\mathsf{T}}_{s}(t)\cdot\mathcal{P}_{\mathrm{safe}}\cdot\boldsymbol{x}_{s}(t)},$ (6c)
where $\mathrm{w}_{\mathrm{com}}$ and $\mathrm{w}_{\mathrm{safe}}$ design parameters. Furthermore, the matrices are
$\mathcal{Q}_{\mathrm{ref}}$ $=\mathrm{diag}\underbrace{\left[{\mathbf{Q}}_{\mathrm{ref}},{\mathbf{Q}}_{\mathrm{ref}},...,{\mathbf{Q}}_{\mathrm{ref}}\right]}_{N\,\mathrm{times}}$
$\mathcal{P}_{\mathrm{safe}}$ $=\mathrm{diag}\underbrace{\left[{\mathbf{P}}_{\mathrm{ref}},{\mathbf{P}}_{\mathrm{ref}},...,{\mathbf{P}}_{\mathrm{ref}}\right]}_{N\,\mathrm{times}}$
in which for the weighting ${\mathbf{Q}}_{\mathrm{ref}}$ and the perturbation ${\mathbf{P}}_{\mathrm{safe}}$ matrices,
${\mathbf{Q}}_{\mathrm{ref}}$ $=\mathrm{diag}\left[0,\mathrm{w}_{\mathrm{ref}_{\mathrm{veh}}},0,\mathrm{w}_{\mathrm{ref}_{\mathrm{ped}}}\right]\,\mathrm{and}$
${\mathbf{P}}_{\mathrm{safe}}$ $=\mathrm{diag}\left[1,0,1,0\right]$
hold. The decision result of the IAMPDM is obtained from the optimization
$\boldsymbol{u}_{s}^{*}$ $=\mathrm{arg}\,\mathrm{min}\,J_{\mathrm{MPC}}\left(\boldsymbol{u}_{s}\right)$ (7a)
$\mathrm{s.t.}\;$ (4) (7b)
$d^{2}_{\mathrm{min}}\leq x_{\mathrm{veh}}^{2}(t)+y_{\mathrm{ped}}^{2}(t)$ (7c)
$0\leq\dot{x}_{\mathrm{veh}}(t)\leq\dot{x}^{\mathrm{max}}_{\mathrm{veh}}$ (7d)
$a_{\mathrm{veh}}^{\mathrm{min}}\leq u(t)\leq a_{\mathrm{veh}}^{\mathrm{max}},$ (7e)
which computed the target acceleration of the AV. The optimization (7) is solved by a Python API of CasADi (Andersson et al., 2019), where the nonlinear programming solver with an interior point optimizer is applied, which can handle both the state (7c), (7d) and input constraints (7e). With these hard constraints, the safety of the IAMPDM can be formally ensured by strictly enforcing operational boundaries.
3.3 Integrating Intention Modeling
The intention of the pedestrian to cross the street is derived from their non-verbal communication, which includes eye contact, gestures, and body posture. These explicit communication signals are referred to as intention. Pedestrians use these cues similarly when interacting with human-driven vehicles: They give way with hand gestures or signal their right of way by looking at the human driver. These explicit communication signals can be determined by machine-learning-based/data-driven detection systems (see e.g. Yang et al. (2022)) and should be taken into account to extend the MPC for the decision-making. Therefore, the crossing intention $I_{\mathrm{ped}}(t)$ of the pedestrian is introduced as a function of time and pedestrian behavior, which can vary between 00 and $1$, which can be used for predicting the actions of the pedestrian. Note that the aim of this paper is not the development of such intention-detection algorithms. It is assumed for this work that they are given since intention-detection algorithms can be found in the literature (Razali et al., 2021), (Rasouli et al., 2022), (Azarmi et al., 2023), (Zhou et al., 2024). To integrate the pedestrian’s crossing intention into the MPC formulation, the following two extensions are implemented:
  • a)
    If the pedestrian is in the safe or near-zone (see Figure 2), then the parameter $\mathrm{w}_{\mathrm{safe}}$ of the MPC is modified based on the crossing intention of the pedestrian
  • b)
    If the velocity of the pedestrian in the safe or near zone is zero, a discount function for the intention is introduced.
The core idea is to use the modified parameters $\mathrm{w}^{*}_{\mathrm{safe}}$ and $d^{*}_{\mathrm{min}}$ in $J_{\mathrm{safe}}$ by introducing
$\mathrm{w}^{*}_{\mathrm{safe}}=\begin{cases}\mathrm{w}_{\mathrm{safe}}\cdot I_{\mathrm{ped}}(t)&\;\text{if pedestrian is not in CZ}\\ \mathrm{w}_{\mathrm{safe}}&\;\mathrm{else}\end{cases}$ (8)
and
$d^{*}_{\mathrm{min}}=\begin{cases}d_{\mathrm{min}}\cdot I_{\mathrm{ped}}(t)&\;\text{if pedestrian is not in CZ}\\ d_{\mathrm{min}}&\;\mathrm{else}.\end{cases}$ (9)
Thus, an IAMPDM is derived, which includes explicit communication elements. This extension introduces a new aspect that hasn’t been applied in state-of-the-art research yet.
However, these elements can sometimes be contradictory, leading to a traffic standstill. Pedestrians often move quickly in the safe zone and slow down in the near zone. They may sometimes wait in the near zone even if they have the right of way or can cross. In such cases, an automated vehicle would stop and wait for the pedestrian [^4] Considering the critical significance of safety, automated vehicles tend to behave conservatively, leading to more frequent stops.. This results in neither the pedestrian nor the vehicle moving, causing a deadlock. As a solution, a discount function of the intention is proposed motivated by game theoretical applications, see e.g. Marín-Solano (2021):
$I^{*}_{\mathrm{ped}}(t)=I_{\mathrm{ped}}(t_{0})\cdot 0.9^{K_{\mathrm{d}}\cdot t},$ (10)
where $K_{\mathrm{d}}$ represents a design parameter, and $t_{0}$ denotes the onset of the interaction between the AV and the pedestrian. Instead of the initial crossing intention of the pedestrian $I_{\mathrm{ped}}(t_{0})$, the discounted value is taken into account.
If either the pedestrian or the vehicle has passed the intersection, no interaction occurs, allowing the vehicle to proceed at its reference speed, which is controlled by a simple velocity-tracking controller.
1 Input: $ped$, $veh$, $I_{\mathrm{ped}}(t_{0})$
2 Output: $veh\_acc$
3
4if $is\_ped\_passed$ or $is\_veh\_passed$ then
5       return velocity_control($ped,veh$)
6else
7       if $is\_discounting\_intention$ then
8             use (10)
9      else
10             use $I_{\mathrm{ped}}(t_{0})$
11
12      (8) and (9)
13       Update vehicle state (4)
14       Update $J_{\mathrm{MPC}}$ (6)
15       $veh\_acc\leftarrow$ Solve (7)
16       return $veh\_acc$
17
Algorithm 1 The IAMPDM Algorithm
Note that the IAMPDM can be handle also. Managing dense pedestrian traffic is indeed a challenge in pedestrian crossing scenarios. In scenarios with many pedestrians crossing simultaneously, the IAMPDM framework can be adapted by interacting with the closest person. Another solution, is that the person with the highest estimated crossing intention is taken into account for the interaction. For both cases, the core working principle of the IAMPDM does not change.
4 Experiment with the Interaction-Aware Model Predictive Decision-Making
In this section, the experiment and its results are presented. The goal of the experiment is to validate the applicability of the IAMPDM and compare it with
  • 1.
    a non-interactive algorithm (NIA) being cautious in interactions with pedestrians and
  • 2.
    a rule-based decision-making (RBDM) introduced in Varga et al. (2023a).
The independent variable of the study design was the selection of the negotiation algorithm (MPDM, RBDM, NIA). Note: The Ethical Commission of WIVW approved the study based on Code of Ethics at WIVW – Summary for project proposals and articles/papers.
4.1 Experimental Simulator Platform
The simulator was set up in an $8\times 4$ meter room to simulate an urban street crossing situation for pedestrians. A screen was placed on one wall of the room, projecting the perspective of a pedestrian. This dynamic view corresponded to that of a pedestrian looking perpendicularly along the street in the crossing direction, see Figure 3. To precisely capture the pedestrian’s position, Vive Trackers were used, with the projected view changing according to the position of the test subjects. The main advantage of this configuration lies in avoiding the motion sickness of a virtual reality system, enabling longer usage and accurately representing natural crossing behavior. In contrast, using a VR headset could cause simulator sickness, and the lack of visual representations of body parts in the virtual environment might lead to unnatural movement behavior. Therefore, we decided against a VR experiment based on these considerations and the advice of the ethical commission. of WIVW.
The experimental setup includes
  • 1
    A simulation computer with SILAB® [^5] SILAB is a software product of WIVW - Würzburg Institute for Traffic Sciences GmbH (www.wivw.de). software framework,
  • 2
    A short-distance projector to display the scenario with the automated vehicle,
  • 3
    A body sensor to track the position and velocity of the test subjects and
  • 4
    An input joystick for detecting the crossing intention of the test subjects.
The advantage of using a joystick as input for the crossing intention helps to overcome the errors of computer vision algorithms that would be necessary to recognize the crossing intention of the pedestrian. Since this work does not attempt to validate such detection algorithms, the use of the joystick is a reasonable solution. Therefore, the experiment is even more rigorous in its objective of comparing different decision-making algorithms in a closed-loop manner.
Figure 3: The simulator setup. Note that the light was turned off during the experiments, and the room was completely dark.
4.2 The Setups of the Decision-Making Algorithms
The critical design parameters of the IAMPDM are shown in Table 1. These parameters facilitate the implementation of either a more aggressive or a more conservative/cautious behavior of the AV.
Parameter Impact of the parameter on
$\mathrm{w}_{\mathrm{safe}}$ Stopping distance
$\mathrm{w}_{\mathrm{com}}$ Deceleration rate
$\mathrm{w}_{\mathrm{ref}_{\mathrm{ped}}}$ Deceleration rate
$\mathrm{w}_{\mathrm{ref}_{\mathrm{veh}}}$ Velocity profiles
$K_{\mathrm{d}}$ Waiting time
Table 1: The most critical parameters and their values of the IAMPDM algorithm
Instead of a manual tuning of these parameters, it is beneficial to apply an automated, optimization-based tuning framework. However, the usage of such automated tuning frameworks from the literature was not suitable for setting up the IAMPDM since they do not take human subjective preferences into account. The inclusion of human preferences into IAMPDM tuning is one challenging aspect since it helps to enhance public acceptance of such technical systems.
An automated tuning framework for pedestrian-AV interaction is presented in Varga et al. (2023a), which utilizes the optimization
$\boldsymbol{\theta}^{*}$ $=\mathrm{arg}\,\mathrm{min}\,J_{\mathrm{glob}}$ (11a)
[…]
to find the optimal parameter vector $\boldsymbol{\theta}$ of an intention-aware decision-making. The global cost function has the form
$J_{\mathrm{glob}}=\int_{t_{0}}^{T_{\mathrm{end}}}k_{1}\cdot t+k_{2}\cdot\left|a^{2}_{\mathrm{max,veh}}(t)\right|-k_{3}\cdot\left|d_{\mathrm{min}}\right|+k_{4}\cdot\frac{1}{TTC(t)}\,\mathrm{d}t,$
from which the optimal parameters of the intention-aware decision-making are obtained. In previous works, the parameters $k_{i},i=\{1,2,3,4\}$ were determined based on optimization goals, such as acceleration profiles and stopping distances. These selections, however, do not account for human preferences. Therefore, for this work, we suggest a semi-automated expert design process to systematically tune the parameters $k_{i}$. The proposed semi-automated expert design of the IAMPDM has the following steps:
Input: $k_{i},i=\{1,2,3,4\}$
Output: $\boldsymbol{\theta}^{*}$
1
2while Overaall result not satisfactory do
3       Setting $k_{i},i=\{1,2,3,4\}$
4       Run (11) with $J_{\mathrm{glob}}$
5       Testing the decision-making
6
Algorithm 2 The algorithm of the iterative design of the intention-aware decision-making including human preferences
Through this iterative design, we can obtain the optimal parameters for IAMPDM integrating human preferences. By including these preferences, we achieve more human-centered optimization results. This human-centered optimization is beneficial because it requires tuning only a few parameters, no matter the algorithm’s complexity, since the parameter vector of IAMPDM consists of
$\boldsymbol{\theta}=\left[\mathrm{w}_{\mathrm{safe}},\mathrm{w}_{\mathrm{com}},\mathrm{w}_{\mathrm{ref}_{\mathrm{ped}}},\mathrm{w}_{\mathrm{ref}_{\mathrm{veh}}},d_{\mathrm{min}},K_{\mathrm{d}},x^{\mathrm{max}}_{\mathrm{veh}},a_{\mathrm{veh}}^{\mathrm{min}},a_{\mathrm{veh}}^{\mathrm{max}}\right]$
that can be difficult to tune manually. Moreover, $k_{i},i={1,2,3,4}$ always preserves their physical meaning, leading to time-efficient human-centered optimization.
The RBDM is also tuned with Algorithm 2; the only difference is that the RBDM has fewer parameters.
To mitigate the risk of injury in urban areas, the non-interactive algorithm assumes that pedestrians near the street edge may cross unpredictably at any time. As a result, the vehicle halts and waits for a duration of $t_{\mathrm{NIA}}$, disregarding pedestrian crossing intentions since it assumes the absence of high-level communication detection. If the pedestrian does not cross, the vehicle starts moving again cautiously after $t_{\mathrm{NIA}}$, in accordance with the IAMPDM discount function, which leads to similar wait times.
4.3 Experiment Design
The three different decision-making algorithms were marked with different colors during the experiment, meaning that they were not revealed to the test subjects. As a result, distinguishing was simplified without using numbers, resulting in fewer matching errors by the test subjects.
Since existing research does not address the human-in-the-decision-loop nature of pedestrian-AV interactions in unsignalised crossings, a key challenge was to instruct test subjects in a way that 1) facilitates interaction but 2) does not result in a fixed sequence of human actions that entirely predetermines their behavior. This balance is challenging, as too much freedom leads to high variability, whereas fully predefined behavior compromises the evaluation of the IAMPDM. Forcing the interaction may result in unrealistic scenarios and misleading results.
This challenge was addressed by instructing the test subjects to mimic altering their decisions: They started to cross the street, but they waited on the roadside. Therefore, we defined four scenarios, which describe approximately how the test subjects should behave, but no fixed sequence of human actions is defined, no, crossing time and distance specifications for specific actions are given in advance. The detailed instructions are given in the Appendix. These scenarios are
  • 1
    Crossing before the vehicle
  • 2
    Remaining and letting the vehicle to cross first
  • 3
    Delayed Crossing before the vehicle
  • 4
    Delayed Remaining and letting the vehicle to cross first.
The experimental procedure was organized as follows: Test subjects started with scenarios 1 and 2, performed in a randomized order. Each scenario involved three crossings using each of the three decision-making algorithms, leading to 9 crossings per scenario. After completing both scenarios, they filled out an intermediate questionnaire to assess the algorithms. This questionnaire was crucial for shaping the test subjects’ perspectives for the final evaluation.
Afterward, the test subjects repeated the procedure with scenarios 3 and 4 in a randomized sequence. They tested all the decision-making algorithms and filled out the intermediate questionnaire again. Finally, they had to answer the final questions, which are used for subjective assessment of the algorithms. Since scenarios 3 and 4 are the most crucial and involve contradictory interactions with the AV, only these were used for objective assessment.
4.4 Objective Goals and Evaluation Criteria
In the following, our null and working hypotheses are presented, and the measures of the experiment are discussed.
The investigation included two hypotheses focusing on objective and subjective measures. For the objective assessment, we have the following null and working hypotheses:
  • H10
    The use of intention-aware decision-making in autonomous vehicles significantly does not alter the crossing times and the criticality of intersection scenarios.
  • H1w
    The use of intention-aware decision-making in autonomous vehicles significantly shortens crossing times and elevates the criticality of intersection scenarios.
The subjective assessment is based on the following null and alternative hypotheses:
  • H20
    The usage of intention-aware decision-making does not lead to a higher user preference in the intersection scenarios.
  • H2w
    The usage of intention-aware decision-making leads to a higher user preference in the intersection scenarios.
To measure our hypotheses, we used three evaluation metrics defined in Westhofen et al. (2023):
  • 1.
    The average Time-to-Collision ($TTC_{\mathrm{avg}}$). $TTC(t)$ is computed such as
    $TTC(t)=\frac{y_{\mathrm{ped}}(t)+x_{\mathrm{veh}}(t)}{\mathrm{max}(v_{\mathrm{veh}}(t),\kappa)},$ (12)
    where $\kappa=0.05$ in order to ensure the numerical stability, in cases of the pedestrian or the vehicle ar at a standstill. From that the average $TTC_{\mathrm{avg}}$ is calculated by
    $TTC_{\mathrm{avg}}=\frac{1}{N}\int_{t_{0}}^{T_{\mathrm{end}}}TTC(t)\,\text{d}t.$ (13)
  • 2.
    Average Deceleration to Safety Time (DST)
    $DST(t)=\frac{1}{2}\frac{v^{2}_{\mathrm{ped}}(t)+v^{2}_{\mathrm{veh}}(t)}{x_{\mathrm{veh}}(t)+y_{\mathrm{ped}}(t)+v_{\mathrm{veh}}(t)\cdot t_{\mathrm{safe}}},$ (14)
    where the safety time is $t_{\mathrm{safe}}=1$. The average $DST_{\mathrm{avg}}$ is calculated by
    $DST_{\mathrm{avg}}=\frac{1}{N}\int_{t_{0}}^{T_{\mathrm{end}}}DST(t)\,\text{d}t.$ (15)
  • 3.
    The completion time of the scenario ($T_{\mathrm{end}}$), defined as the time after the vehicle or pedestrian has left the collision zone.
Besides these objective metrics, the test subjects had to assess the decision-making algorithm in the final questions to obtain their subject impressions. These are
Question 1:
Which decision-making did you find overall to be the most appropriate concerning the experienced situations?
Question 2:
Please now evaluate the decision-making algorithms based on your experiences in all the situations: Can you imagine a red/green/blue vehicle with this decision-making algorithm operating in real traffic?
What do you think of red/green/blue decision-making in general?
The possible answers are given in Table 2.
not
at all
very
little
little medium strong
very
strong
0 1,2,3 4,5,6 7,8,9 10,11,12 13,14,15
Table 2: Answer Table for the test-subjects
Note hypothesis H1 is evaluated by the objective metrics $TTC$, $DST$ and $T_{\mathrm{end}}$. For the assessment of hypothesis H2 the final questions are used.
5 Results and Discussion
This section presents the results of our experiment and provides an in-depth discussion. Please note that the Delayed Crossing and Delayed Remaining scenarios cause different overall behaviors in the pedestrian and the automated vehicle, making it necessary to assess them separately. For both objective and subjective results, the outliers are removed using the inter-quartile range method, see e.g. Dash et al. (2023).
5.1 Objective Results
The resulting average values of the measures $TTC_{\mathrm{avg}}$, $DST_{\mathrm{avg}}$ and $T_{\mathrm{end}}$ with their standard deviations are presented in Table 3 for the Delayed Crossing case and in Table 4 for the Delayed Remaining case. Furthermore, Figure 4 and Figure 5 show the box plots of the objective results.
It can be seen that the NIA has the largest $T_{\mathrm{end}}$ for both scenarios. On the other hand, the IAMPDM and RBDM have comparable shorter crossing times. This means that negotiations between a pedestrian and an AV took shorter, indicating that using intention-aware decision-making algorithms can lead to smoother traffic.
On the other hand, the criticality metrics (DST and TTC) show that the scenarios became more critical: The TTC was smaller for both IAMPDM and RBDM compared to NIA. Furthermore, the DST values were smaller if the NIA was used [^6] Note that the smaller the DST, the less critical the scenario. On the other hand, larger TTC values mean safer (less critical) scenarios..
To assess whether these differences between the decision-making algorithms are statistically significant, statistical tests are performed. First, the Kruskal-Wallis test is utilized for evaluating H1, because all three samples are compared, see e.g. Dalgaard (2008). The degrees of freedom of this test are $df=2$, and the significance level is chosen to $\alpha=0.01.$ Its null hypothesis is that there is no difference between the three decision-making algorithms. This hypothesis is declined if $\mathcal{H}\geq\mathcal{X}_{df,\alpha}^{2}$ holds, where $\mathcal{X}^{2}_{df=2,\alpha=0.01}=9.21$.
$T_{\mathrm{end}}$ in $s$ $TTC_{\mathrm{avg}}$ in $s$ $DST_{\mathrm{avg}}$ in $\frac{m}{s^{2}}$
$\mu_{\mathrm{NIA}}$ 28.49 42.09 2.02
$\sigma_{\mathrm{NIA}}$ 3.47 36.75 1.33
$\mu_{\mathrm{RBDM}}$ 18.87 7.67 2.69
$\sigma_{\mathrm{RBDM}}$ 5.12 0.56 1.16
$\mu_{\mathrm{IAMPDM}}$ 27.51 7.91 3.01
$\sigma_{\mathrm{IAMPDM}}$ 17.63 1.43 0.92
Table 3: Mean values and their standard deviation of the objective metrics in case of Delayed Crossing
$T_{\mathrm{end}}$ in $s$ $TTC_{\mathrm{avg}}$ in $s$ $DST_{\mathrm{avg}}$ in $\frac{m}{s^{2}}$
$\mu_{\mathrm{NIA}}$ 29.71 41.27 1.85
$\sigma_{\mathrm{NIA}}$ 1.65 35.72 1.38
$\mu_{\mathrm{RBDM}}$ 16.12 7.92 3.01
$\sigma_{\mathrm{RBDM}}$ 2.56 0.17 0.57
$\mu_{\mathrm{IAMPDM}}$ 17.18 7.59 2.71
$\sigma_{\mathrm{IAMPDM}}$ 4.09 0.10 0.37
Table 4: Mean values and their standard deviation of the objective metrics in case of Delayed Remaining
In case of Delayed Crossing, the following $\mathcal{H}$ values are obtained
$\mathcal{H}_{Crs_{T_{\mathrm{end}}}}$ $=16.64$
$\mathcal{H}_{Crs_{DST}}$ $=6.82$
$\mathcal{H}_{Crs_{TTC}}$ $=22.42.$
Since $\mathcal{H}_{i}\geq\mathcal{X}_{df,\alpha}^{2}$ hold for $i=\{Crs_{T_{\mathrm{end}},Crs_{TTC}}\}$, we can reject the null hypothesis H10 for the Delayed Crossing scenarios that the average $T_{\mathrm{end}}$ and $TTC$ are same for all three decision-making. This indicates that the choice of the decision-making algorithms leads to statistically significant differences in $T_{\mathrm{end}}$ and $TTC$. On the other hand, $\mathcal{H}_{i}\geq\mathcal{X}_{df,\alpha}^{2}$ does not hold for $i=Crs_{DST}$ meaning the we cannot reject the null hypothesis of H1 in case of $DST$. In case of Delayed Remaining, the following $\mathcal{H}$ values are obtained
$\mathcal{H}_{Rem_{T_{\mathrm{end}}}}$ $=32.44$
$\mathcal{H}_{Rem_{DST}}$ $=9.91$
$\mathcal{H}_{Rem_{TTC}}$ $=30.09$
Since $\mathcal{H}_{i}\geq\mathcal{X}^{2}_{df,\alpha}$ is satisfied for $i=\{Rem_{T_{\mathrm{end}}},Rem_{DST},Rem_{TTC}\}$, we can conclude that the choice of decision-making algorithms has statistically significant effects on $T_{\mathrm{end}}$, $DST$ and $TTC$. In addition, Mann–Whitney–Wilcoxon tests are carried out to compare IAMPDM and RBDM. As it can be seen from Table 5, aside from $TCC$ in Delayed Remaining case, there is no statistical difference between IAMPDM and RBDM.
(a) Average crossing times
(b) Average DSTs
(c) Average TTCs for Delayed Crossing
Figure 4: Results in case of Delayed Crossing
(a) Average crossing times
(b) Average DSTs
(c) Average TTCs for Delayed Crossing
Figure 5: Results in case of Delayed Remaining
Parameter $p$-value
$p_{\mathrm{Crs_{T_{\mathrm{end}}}}}$ $0.34$
$p_{\mathrm{Crs_{DST}}}$ $0.45$
$p_{\mathrm{Crs_{TTC}}}$ $0.51$
$p_{\mathrm{Rem_{T_{\mathrm{end}}}}}$ $0.73$
$p_{\mathrm{Rem_{DST}}}$ $0.08$
[…]
Table 5: The resulting $p$-values of Mann–Whitney–Wilcoxon tests to compare IAMPDM and RBDM
5.2 Subjective Results
For the testing of H2, the results of the final questions are analyzed, for which the test subject Nr. 21 is excluded since the person could not recall which decision-making was which. First, the preferences of the test subject are given in Table 6. It can be seen that most of the test subjects have chosen intention-aware decision-making. The results of Q2 from the final questions are given in Table 7. The raw data of Q2 is included in B. It can be seen that in the subjective assessment, both RBDM and the IAMPDM possess larger mean values than NIA. The Kruskal-Wallis test is conducted to test the difference for statistical significance. The result is
$\mathcal{H}_{\mathrm{Subj}}=14.56,$
which indicates that the choice between intention-aware and non-interactive decision-making algorithms leads to statistically significant differences in the preference of the human test subjects. For a comparison between the IAMPDM and RBDM, a Mann–Whitney–Wilcoxon test is carried out, which yields
$p_{\mathrm{Subj}}=0.154,$
indicating that the difference between IAMPDM and RBDM is not significant.
Preferences NIA 5 RBDM 14 IAMPDM 5 Table 6: Preferences of the test subjects based on Q1 Final Question Score $\mu_{\mathrm{NIA}}$ 7.04 $\sigma_{\mathrm{NIA}}$ 3.629 $\mu_{\mathrm{RBDM}}$ 11.04 $\sigma_{\mathrm{RBDM}}$ 3.665 $\mu_{\mathrm{IAMPDM}}$ 10.00 $\sigma_{\mathrm{IAMPDM}}$ 2.690 Table 7: Mean values and standard deviations of Q2
Figure 6: Results of the subjective assessment
5.3 Discussion and Limitations
One of the limitations of our experimental setup lies in the assumption of perfect crossing intention detection via joystick input. While this simplified approach enabled controlled testing, real-world intention recognition involves inherent uncertainties due to sensor noise, occlusions, and ambiguous pedestrian behavior.
It is necessary to highlight that higher TTC values are associated with more natural driving behavior of the AV. This preference is reflected in the test subjects choosing the RBDM and IAMPDM over the more conservative NIA algorithm. Additionally, the test subjects provided insights into the general question What do you think of red/green/blue decision-making in general?, which enhanced our understanding of the proposed algorithms. From the study results, we can draw interesting conclusions. A closer examination of the objective results reveals that the intention-aware decision-making algorithms consistently outperform the NIA in all aspects. However, there is no significant difference between IAMPDM and RB, except in the average TTCs concerning Delayed Remaining (see Table 5). Additionally, we observed sporadic instances of extremely high crossing times for IAMPDM in Figure 5(a). These outliers primarily occurred when pedestrians hesitated mid-crossing, triggering repeated velocity adjustments of the AV. This reactive behavior suggests the need for better anticipation of pedestrian motion variability during the prediction.
For the other cases and measures, no significant differences between IAMPDM and RB were observed in this study. These results suggest that a simple algorithm can achieve outcomes comparable to more complex methods in simple urban traffic scenarios. This implies that distinguishing between simple scenarios and complex scenarios is necessary for the development and validation of interaction-aware algorithms for autonomous vehicles. A simple scenario is defined as dyadic interactions with single pedestrians in controlled crossing situations, characterized by predictable trajectories, as in our experiment setup. Complex scenarios involve multiple interacting agents, mixed traffic modes, and ambiguous navigation goals. This distinction is crucial because while rule-based systems suffice for simple cases, complex scenarios demand the adaptive reasoning capabilities of model-based approaches like IAMPDM. The model (1) is also designed for more complex scenarios. It can extend to situations with multiple interacting agents. This flexibility can make it more applicable to a broader range of pedestrian-vehicle interactions, including those involving dynamic or unpredictable behaviors.
Moreover, the subjective results offer interesting insights from the human-in-the-decision-loop experiment. The test subjects, who preferred the NIA, indicated that a more conservative approach could increase safety and be more favored by pedestrians. They had the comment such
  • 1.
    Too passive and slow for my taste, but that’s not wrong in normal traffic. Then misunderstandings are not so tragic.
  • 2.
    As a pedestrian, I had the feeling that it [the vehicle] looked out for me in every situation and always gave me the opportunity to cross the road. Overall, I felt safest with [NIA].
  • 3.
    Decision-making for the pedestrians pleasant.
Most of the test subjects liked the RBDM more and criticized the long waiting time of the NIA. An interesting result is that the RBDM leads to better subjective results compared to IAMPDM: Most of the test subjects preferred RBDM over the IAMPDM, see Table 6. However, there is no significant difference in the objective results; see Table 7. Furthermore, most of them were not able not distinguish between RBDM and IAMPDM. They left comments like
  • 1.
    I found hard to the difference between [RBDM] and [IAMPDM], but [RBDM] seemed a bit more thoughtful. Reacted better and more appropriately than [IAMPDM].
  • 2.
    [In case of IAMPDM:] To [RBDM] no difference detectable
  • 3.
    [In case of RBDM:] no big difference to [IAMPDM]
The reason for this result is the tuning framework, since the human factors component was tuned for both RBDM and IAMPDM in a similar manner, consequently, the subjective results were also similar. Some general comments were that the test subjects were not sure whether the vehicle recognized them or not. To overcome this, implementing a human-machine interface (HMI) for a vehicle to communicate with pedestrians could significantly enhance trust. When a pedestrian sees a clear signal from the vehicle indicating it has detected them, it reduces uncertainty and potential accidents.
The findings suggest that simply increasing the complexity of intention-aware decision-making algorithms does not necessarily lead to better performance or greater pedestrian preference. Additionally, the conditions under which these more complex and reliable model-based decision-making algorithms can achieve a safe scenario need to be validated in further studies.
6 Summary
This paper proposes an interaction-aware model predictive decision-making system and its real-time implementation for automated vehicles interacting with pedestrians in urban, low-speed scenarios. Furthermore, we designed a human-in-the-decision-loop study to examine the human-automation action-reaction cycle. The human-in-the-decision-loop study, conducted with 25 participants, demonstrated that intention-aware decision-making algorithms lead to faster resolution of human-automation negotiations and improved subjective evaluations. These findings highlight the significance of intention-aware decision-making for autonomous vehicles in urban settings. Our future work will feature an adaptive decision-making algorithm based on Varga (2024). Furthermore, we plan to integrate a probabilistic estimation of human crossing intentions by modeling their stochastic behavior, see Škugor et al. (2023).
CRediT authorship contribution statement
Balint Varga: Writing – original draft, Writing – review & editing, Conceptualization, Formal analysis, Methodology, Software, Visualization, Validation. Thomas Brand: Writing – review & editing, Conceptualization, Data curation, Methodology, Software. Marcus Schmitz: Writing – review & editing, Methodology, Conceptualization. Ehsan Hashemi: Writing – review & editing, Investigation, Conceptualization.
Acknowledgments
This work was supported by the Federal Ministry for Economic Affairs and Climate Action, in the New Vehicle and System Technologies research initiative with Project number 19A21008D.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix A Instruction Materials
Crossing
“Please proceed to the markings. In this situation, you want to cross the road in front of the vehicle. Behave as you would with a real vehicle and driver. Please only interact with the vehicle when you think the vehicle is aware of you. Remember to press the button when you want to signal that you want to cross the road. Keep the button pressed as long as you want to cross the road in front of the vehicle. Please remember that you should only cross the road if you feel safe, otherwise do not cross the road.
Remain
“Please proceed to the markings. In this situation, you only want to cross the road after the vehicle. Behave as you would with a real vehicle and driver. Remember to press the button when you want to signal that you want to cross the road. Keep the button pressed as long as you want to cross the road in front of the vehicle.”
Delayed crossing
“Please proceed to the markings. In this situation, you actually want to cross the road after the vehicle. However, you change your mind and spontaneously decide to cross the road before the vehicle. Behave as you would with a real vehicle and driver. Please only interact with the vehicle when you think the vehicle is aware of you. Remember to press the button when you want to signal that you want to cross the road. Keep the button pressed as long as you want to cross the road in front of the vehicle. Please remember that you should only cross the road if you feel safe, otherwise do not cross the road.”
Delayed stay
“Please move to the markings. In this situation, you actually want to cross the road in front of the vehicle. However, you change your mind and spontaneously decide to cross the road after the vehicle. Behave as you would with a real vehicle and driver. Remember to press the button when you want to signal that you want to cross the road. Keep the button pressed as long as you want to cross the road in front of the vehicle.”
Appendix B Subjective Results
IAMPDM RBDM NIA
1 8 8 8
2 1 14 1
3 11 9 6
4 11 13 9
5 9 14 0
6 10 8 4
7 7 11 7
8 6 12 8
9 14 12 6
10 9 12 1
11 7 6 10
12 14 15 10
13 12 15 4
14 9 9 4
15 10 10 5
16 11 15 9
17 9 5 13
18 10 14 11
19 10 13 10
20 4 8 12
21 - - -
22 11 12 7
23 15 15 8
24 13 14 4
25 1 1 12
Table 8: The subjective assessment for IAMPDM, RBDM, and NIA

### v2

@@
 Abstract
-This paper presents the experimental validation of an interaction-aware model predictive decision-making (IAMPDM) approach in the course of a simulator study.
-The proposed IAMPDM uses a model of the pedestrian, which simultaneously predicts their future trajectories and characterizes the interaction between the pedestrian and the automated vehicle.
-The main benefit of the proposed concept and the experiment is that the interaction between the pedestrian and the socially compliant autonomous vehicle leads to smoother traffic.
-Furthermore, the experiment features a novel human-in-the-decision-loop aspect, meaning that the test subjects have no expected behavior or defined sequence of their actions, better imitating real traffic scenarios.
-Results show that intention-aware decision-making algorithms are more effective in realistic conditions and contribute to smoother traffic flow than state-of-the-art solutions.
-Furthermore, the findings emphasize the crucial impact of intention-aware decision-making on autonomous vehicle performance in urban areas and the need for further research.
+Autonomous vehicles must negotiate with pedestrians in ways that are not only safe, but also legible and socially acceptable.
+This paper presents an interaction-aware model predictive decision-making (IAMPDM) framework that couples a gap-acceptance-inspired pedestrian intention model with model predictive control (MPC) to coordinate intent reasoning and vehicle control in real time.
+The pedestrian model maps time-to-collision (TTC) to a continuous crossing-likelihood and incorporates the discounting of the pedestrian’s crossing intention to represent hesitation and potential standstills.
+The resulting signals then scale MPC safety costs and minimum-distance constraints.
+We implement IAMPDM in a projection-based, motion-tracked simulator and compare it with a rule-based intention-aware controller (RBDM) and a conservative non-interactive baseline (NIA).
+In a human-in-the-decision-loop study with 25 participants (one excluded from subjective analyses), intention-aware methods generally reduced scenario completion time relative to NIA in the tested scenarios, while also producing tighter surrogate safety margins (e.g., TTC- and distance-based measures).
+Across most metrics, IAMPDM and RBDM were not statistically distinguishable, except for TTC in one scenario.
+Participants’ subjective ratings tended to favor the intention-aware approaches over the non-interactive baseline.
+Finally, we translate these simulator observations into concrete tuning guidance.
 keywords:
@@
 1 Introduction
-Highly automated and autonomous vehicles (AVs) are increasingly becoming a part of our daily lives (Elliott et al., 2019)(Zhou et al., 2022).
-Integrating these systems into society largely depends on the trust of vulnerable road users, such as cyclists and pedestrians, whose safety is essential (Bouzidi and Hashemi, 2023).
+Highly automated and autonomous vehicles (AVs) are increasingly becoming a part of our daily lives Zhou et al.
+[2022].
+Integrating these systems into society largely depends on the trust of vulnerable road users, such as cyclists and pedestrians, whose safety is essential Bouzidi and Hashemi [2023].
 Incidents involving automated driving functions often make headlines, fostering public skepticism towards these technologies.
-Consequently, significant research efforts are directed at enhancing automated vehicles with advanced communication channels and decision-making algorithms designed to manage complex scenarios.
+Consequently, significant research efforts are directed at enhancing automated vehicles with advanced communication channels and decision-making algorithms designed to manage complex interaction scenarios, see Markkula et al.
+[2020a].
 For example, in urban environments and city centers – where vehicles move at slower speeds, and pedestrians may cross unexpectedly – an effective human-machine interaction is critical to fostering trust in these systems.
-Figure 1: The pedestrian crosses the street at an unsignalised intersection interacting with an autonomous vehicle. With the courtesy of version1 GmbH.
-In recent years, advanced algorithms have enabled automated vehicles to interact with pedestrians and cyclists (Hu et al., 2018), (Razmi Rad et al., 2020).
-The efficiency of these algorithms from the literature is often analyzed either in pure simulation setups (Kampitakis et al., 2023) or in human-in-the-loop tests see e.
-g.
-Brand et al.
-(2024).
-These human-in-the-loop experiments [^1] Note that the terms experiment and study are used interchangeably in the course of this paper.
-are usually conducted in virtual environments (Razmi Rad et al., 2020) (Luu et al., 2022).
-However, literature-based experiments restrict decision possibilities: the vehicle’s behaviors or the test subjects’ expected reactions are predefined, making the experiment more controlled and analysis more accessible, limiting their application in real-world scenarios.
-Therefore, this work validates an interaction-aware model predictive decision-making (IAMPDM) within a simulator experiment using a human-in-the-decision-loop setup, where test subjects’ reactions are unspecified to replicate real-world scenarios better.
-Since field studies with actual vehicles can be risky, human-in-the-loop simulations offer a safer alternative, see e.g Razmi Rad et al.
-(2020), Tran et al.
-(2021), Bindschädel et al.
-(2021) or Kalantari et al.
-(2023).
-This paper provides a human-in-the-decision-loop simulator study for the first time with the following contributions:
+Figure 1 illustrates a typical urban scenario where a pedestrian crosses the street at an unsignalized intersection while interacting with an autonomous vehicle.
+Figure 1: The pedestrian crosses the street at an unsignalized intersection interacting with an autonomous vehicle. With the courtesy of version1 GmbH
+In recent years, advances in machine learning and control – especially game-theoretic models and reinforcement learning (RL) – have improved automated vehicles’ ability to interact with pedestrians and cyclists by capturing the negotiation and decision dynamics that arise in mixed-traffic situations, Hu et al.
+[2018], Razmi Rad et al.
+[2020], Kampitakis et al.
+[2023], Kalantari et al.
+[2023] The performance and safety of these interaction strategies reported in the literature are commonly evaluated under two principal paradigms.
+The first is pure simulation - large-scale, repeatable virtual environments that allow researchers to explore many scenarios and parameter settings efficiently Luu et al.
+[2022], Kampitakis et al.
+[2023].
+The second paradigm embeds humans into the evaluation loop, either through human-in-the-loop studies or controlled user experiments, which can reveal important aspects of human behavior, trust, and perceived safety that simulations alone cannot capture Tran et al.
+[2021], Bindschädel et al.
+[2021].
+While both approaches are indispensable, literature-based experiments typically impose constraints on the decision space: vehicle actions (e.g., braking profiles, trajectory choices) and the range of participants’ responses are often predefined to make analysis tractable and results reproducible.
+This strong experimental control facilitates statistical comparison across algorithms and conditions but comes at the cost of external validity - the restricted set of behaviors and responses can fail to capture the unpredictability and heterogeneity of real-world interactions.
+As a result, findings derived from such controlled setups may overestimate robustness or underrepresent rare but safety-critical behaviors, limiting the applicability of conclusions when algorithms are deployed outside laboratory conditions.
+Therefore, in this work, we address those limitations by proposing and validating an interaction aware model predictive decision-making (IAMPDM) approach in a simulator using a true human-in-the-decision-loop setup where participants are free to choose their actions.
+Simulator studies are a practical and ethically safer alternative to on-road trials for examining unconstrained human–AV negotiation Tran et al.
+[2021].
+Building on this rationale, the present paper reports a simulator experiment that preserves participant freedom while enabling a systematic comparison of decision-making strategies.
+The main contributions are:
   • 1.
-    Adapting the IAMPDM for the real-time experiments,
+    We adapt the IAMPDM for real-time experiments,
   • 2.
-    Designing a suitable experimental human-in-the-decision-loop simulator study, and
+    We design a novel human-in-the-decision-loop simulator study that balances naturalistic validity and experimental control.
   • 3.
-    Conducting the experiment with 25 participants, offering analysis and recommendations for the real-world implementation of the proposed IAMPDM.
+    Finally, we conduct an experiment with 25 participants, and analyze the findings to derive recommendations for real - world implementation of IAMPDM.
 The paper is structured as follows: In Section 2, the state of the art of intention-aware automated vehicles is presented.
@@
 2 State of the Art
-This section provides a short comprehensive literature review of interaction-aware decision-making algorithms, which became large in recent years, since the topic is getting more and more relevant.
-Therefore, the following state-of-the-art overview focuses on
-  • 1.
-    decision-making methods, models and algorithms [^2] For simplicity, the terms decision-making methods, models, and algorithms will collectively be referred to as decision-making concepts. and
-  • 2.
-    experiments and studies, which can verify and validate these concepts.
-For more detailed reviews, the reader is referred to Predhumeau et al. (2023).
+The upcoming sections review key conceptual frameworks, contrasts model-free and model-based approaches, and discusses validation methods for autonomous vehicle decision-making algorithms, with emphasis on interaction-aware systems.
+2.1 Conceptual Frameworks for Interaction
+Defining “interaction” in the context of mixed human-  automated traffic is a central challenge for the development of advanced driver-assistance and autonomous vehicle systems.
+According to the conceptual framework presented by Markkula et al.
+[2020a], interaction encompasses any situation in which the actions of one road user are shaped by the observed or anticipated behavior of another.
+This includes not only direct collision avoidance but also the exchange of information, either implicitly through kinematic cues (such as deceleration profiles or time-to-collision), or explicitly via signals and gestures.
+The framework of Markkula et al.
+[2020a] distinguishes between implicit communication - for example, the yielding intent communicated by gradual deceleration - and explicit communication, such as eye contact or hand gestures.
+The literature increasingly emphasizes the importance of both, as urban traffic negotiation is a continuous, communicative process involving prediction, negotiation, and mutual adaptation.
+Recent surveys, Wang et al.
+[2022] and Crosato et al.
+[2024], catalogue the broad range of computational models developed for social interaction in AVs.
+These reviews cover approaches from game theory, which model strategic reasoning between agents, to frameworks grounded in social psychology, which seek to explain and simulate human-like negotiation, cooperation, and norm adherence in traffic.
 2.2 Decision-Making Concepts
-In the literature, decision-making concepts can be categorized into two main groups: model-based and model-free approaches.
-2.1.1 Model-free Approaches
-Model-free concepts do not utilize physics-based or logic-based models.
-Instead, they learn optimal behaviors through trial and error by directly interacting with the environment.
-These methods use environmental feedback to refine decision-making, often employing algorithms such as Q-learning or policy gradient methods.
-Generally, model-free approaches are well-suited for complex scenarios, such as modeling human behavior.
-A review on this topic can be found in Di and Shi (2021).
+2.2.1 Model-Free, Learning-Based Approaches
+Model-free approaches do not rely on physics-based or logic-based models.
+Instead, they learn optimal behaviors through trial and error by directly interacting with the environment, embodying an end-to-end learning paradigm.
+These methods leverage environmental feedback to refine decision-making policies, often employing algorithms such as Q-learning or policy gradient methods.
+A comprehensive review of this topic can be found in Di and Shi [2021].
 In Russo et al.
-(2021), a Reinforcement Learning (RL) approach is presented addressing pedestrian collision avoidance in autonomous driving systems.
-The focus is on managing unexpected pedestrian crossings and tracking a specific trajectory.
-The Deep Deterministic Policy Gradient (DDPG) algorithm is utilized to learn continuous actions.
-The paper discusses the reward function, agent architecture, and environment model used for training and testing the DDPG-based agent through numerical simulations.
-Similar, data-based concepts are presented in Raghuram Kadali et al.
-(2014), Rezaee et al.
-(2021), Trumpp et al.
-(2022) and Zhang et al.
-(2022).
-However, the main drawbacks of these model-free concepts are that they cannot easily ensure safety or guarantee the reliability needed in high-risk environments, such as AV-pedestrian interactions.
-This limitation means that pure learning-based systems are not mature enough to be safely used in real traffic situations.
-It is to be expected that model-free decision-making approaches will become easier to validate and verify for traffic admission, for which one typical method is the combination of model-free and model-based elements as discussed in Korbmacher and Tordeux (2022) or Derajić et al.
-(2024).
-2.1.2 Model-based Approaches
-On the other hand, model-based concepts utilize mathematical models to describe AV-pedestrian interactions using logical or physical principles.
-These methods rely on this predefined model to predict future states and make decisions.
+[2021], an RL approach is presented that addresses pedestrian collision avoidance in autonomous driving systems.
+The focus lies on managing unexpected pedestrian crossings while maintaining trajectory tracking performance.
+The Deep Deterministic Policy Gradient (DDPG) algorithm is utilized to learn continuous control actions in a high-dimensional state space.
+The authors discuss the reward function design, agent architecture, and environment model used for training and testing the DDPG-based agent through numerical simulations.
+Similar data-driven concepts are presented in Raghuram Kadali et al.
+[2014], Rezaee et al.
+[2021], Trumpp et al.
+[2022] and Zhang et al.
+[2022], demonstrating the breadth of learning-based approaches to pedestrian interaction modeling.
+Additional recent approaches have incorporated social value orientation to shape interaction policies, enabling negotiation styles from assertive to cooperative Crosato et al.
+[2022].
+While effective at modeling subtle dynamics, these methods remain data-intensive and challenging to verify for safety.
+However, the main drawback of these model-free concepts is that they cannot easily ensure safety or provide formal guarantees on the reliability required in high-risk environments such as AV-pedestrian interactions.
+The black-box nature of learned policies makes verification challenging, and their performance may degrade in scenarios not adequately represented in training data.
+This limitation means that pure learning-based systems are not yet mature enough to be safely deployed in real traffic situations without additional safeguards.
+It is expected that model-free decision-making approaches will become easier to validate and verify for traffic admission through the combination of model-free and model-based elements, as discussed in Korbmacher and Tordeux [2022] and Derajić et al.
+[2024].
+Such hybrid architectures aim to leverage the adaptability of learning-based methods while retaining the interpretability and safety guarantees of model-based approaches.
+2.2.2 Mechanistic Model-Based Approaches
+In contrast, mechanistic model-based approaches commonly employ analytic criteria such as gap acceptance and time-to-collision (TTC) thresholds.
+While these criteria are more straightforward to implement and provide a baseline for decision-making, a growing body of research demonstrates that human crossing behavior is richer and depends on a combination of cues.
+For example, Tian et al.
+[2023] found that pedestrians rely not just on TTC, but also on the detailed parameters of vehicle deceleration when assessing whether it is safe to cross.
+In Domeyer et al.
+[2022], a perceptual model is established based on the geometrical relations between drivers and pedestrians, emphasizing the implications for vehicle automation.
+It explores the coupling of driver and pedestrian behavior.
+However, only a simulation analysis is provided, with no strong indications for real-world usage.
+Game theoretical models are presented in Ezzati Amini et al.
+[2024], Skugor et al.
+[2020] and Pavelko et al.
+[2022].
+In these works, the interaction between the pedestrian and vehicle is modeled as a game with two players who optimize their own objective functions through their joint actions.
+These models can represent the interaction between the two players, but they are not suitable for predicting their future joint actions.
+On the other hand, in Chen et al.
+[2023], a model predictive control (MPC) formulation is proposed, which can predict the future behavior of the pedestrian.
+In Škugor et al.
+[2023], the model predictive formulation is extended with a stochastic component, providing a more realistic overall behavior of the proposed algorithm.
+More recently, perceptually-grounded predictive models have been proposed that explicitly account for how pedestrians integrate deceleration cues and gap evolution when forming crossing decisions Tian et al.
+[2025a], and probabilistic vehicle–pedestrian interaction models have been developed to plan optimal implicit-communication behaviors under intent uncertainty Amann et al.
+[2025].
+Building on these findings, perceptually plausible models have been introduced that integrate multiple sensory cues and contextual information into a continuous decision process.
+These methods rely on a predefined model to predict future states and make decisions.
 Furthermore, they can handle constraints and uncertainties in the system-design phase, leading to improved safety and reliability.
 More detailed overviews of model-based methods can be found in Camara et al.
-(2021), Gulzar et al.
-(2021) or Predhumeau et al.
-(2023).
-In Domeyer et al.
-(2022), a perceptual model is established based on the geometrical relations between drivers and pedestrians, emphasizing the implications for vehicle automation.
-It explores the coupling of driver and pedestrian.
-However, only a simulation analysis is provided, with no strong indications for real-world usage.
-Game theoretical models are presented in Cheng et al.
-(2015), Skugor et al.
-(2020) and Pavelko et al.
-(2022).
-In these works, the interaction between the pedestrian and vehicle is modeled as a game with two players who optimize their own objective functions by their joint actions.
-These models can model the interaction between the two players, but they are not suitable to predict their future joint actions.
-On the other hand, in Chen et al.
-(2023), a model predictive control (MPC) formulation is proposed, which can predict the future behavior of the pedestrian.
-In Škugor et al.
-(2023), the model predictive formulation is extended with a stochastic component, providing a more realistic overall behavior of the proposed algorithm.
-However, these concepts often have a high computational complexity, therefore they are often not suitable for real-time application on a state-of-the-art automotive hardware.
-2.2 Validation of Decision-Making Concepts
-Since the main contribution of the work is the experimental study with a human-in-the-decision-loop character, the aftercoming part of the overview focuses on pedestrian-autonomous vehicle experiments from the literature.
-Reviews on the virtual reality studies of pedestrian-automated vehicle interaction can be found in Tran et al.
-(2021) Zhou et al.
-(2022).
-The two main focuses in the literature are the impact of the external human-machine interface (eHMI) on trust and safety.
-For instance, in Prattico et al.
-(2021), various eHMI concepts are compared, and the influence of the eHMI on the test subjects’ trust is investigated.
-Similarly, in Feng et al.
-(2023), the impact of eHMI on pedestrian behavior at road crossings is analyzed in a virtual reality setting.
-The variation of the communication strategies of the automated vehicle is analyzed in Löcken et al.
-(2019).
-The authors proposed five groups for the different communication categories, which are implemented and tested in a virtual reality setup.
-The results showed that the higher the complexity, the more the test subjects trusted the technical system.
-The study in Yang et al.
-(2024) focuses on how pedestrians’ head movements should be interpreted in the presence of automated vehicles at a virtual crossroad, offering insights into non-verbal cues and communication patterns in pedestrian-vehicle interaction for enhanced safety.
-However, the test subjects in this study are not in the decision loop, which means that the interaction between humans and automation has a negligible effect on the study results.
-Thus, the works from the state of the art provide a wide range of experiments and investigations, however, none of these focus on the effect of various decision-making algorithms on the preference of the pedestrians.
+[2021], Gulzar et al.
+[2021] or Predhumeau et al.
+[2023].
+2.3 Social Compliance and Empirical Validation
+Interactions in traffic are multi-agent processes that combine perception, implicit signaling, and negotiation.
+Social compliance (intelligibility, predictability, comfort) is crucial for acceptance: at uncontrolled crossings AVs face a trade-off between conservative and aggressive behaviors, so decision policies must balance physical safety with legibility and social objectives Tian et al.
+[2025b].
+VR studies and human–machine interfaces (eHMI) research (text, icons, lights, kinematics) have probed trust, workload, and communication effectiveness, and work on implicit cues (e.g., head movements) shows how pedestrians infer intent Zhou et al.
+[2022], Löcken et al.
+[2019], Yang et al.
+[2024].
+Nevertheless, many experiments constrain participant influence on AV policies, limiting mutual adaptation and leaving few studies that directly compare different decision-making algorithms under truly interactive, human-in-the-loop conditions.
 2.4 Shortcoming of the State-of-the-Art Methods
-The decision-making and control algorithms from the state of the art have shortcomings that 1) they are either too complex for real-time implementation on general automotive hardware or 2) they are hard to validate in order to pass motor vehicle registration of AVs.
-On the other hand, experimental studies from the literature usually use predefined strategies for autonomous vehicles in the case of human-in-the-loop investigations.
-However, human-in-the-decision-making-loop is an inevitable extension for the further analysis of human factors.
-This means that the experiments should have no predefined behavior, which can have an impact on the results of the decision-making.
-Therefore, our work addresses these challenges and provides an IAMPDM that can run in real time.
-Furthermore, this work also provides an investigation having a human-in-the-decision-making-loop character.
-3 Adjustment of the Model Predictive Algorithm
+Despite impressive progress, current state-of-the-art methods face several limitations.
+Model-free approaches often lack interpretability and formal safety guarantees, hindering certification.
+Model-based approaches can become computationally heavy and may not scale to dense, multi-agent scenarios, and the decision-making and control algorithms from the state of the art are therefore often too complex for real-time implementation on general automotive hardware or hard to validate for motor vehicle registration.
+Moreover, many experiments rely on scripted behaviors or do not put humans fully into the decision loop, limiting ecological validity — human in the decision making loop experiments with no predefined behavior are an inevitable extension for human factors analysis.
+Finally, algorithms rarely consider social compliance explicitly, and there is a paucity of work on quantifying safety efficiency trade-offs.
+Therefore, our work addresses these challenges and provides an IAMPDM that can run in real time; furthermore, this work also provides an investigation having a human‑in‑the‑decision‑making‑loop character.
+3 Improvements of the Model Predictive Decision Making Algorithm
 3.1 Model Description
@@
 The core idea of this IAMPDM is presented in Varga et al.
-(2023b) that we used for our implementation; and we validated the model using the data from Yang et al.
-(2019).
-The proposed model uses a sigmoid function based on the time-to-collision (TTC) values to model the probability of the pedestrian’s velocity changes.
-In the following, a detailed and more computationally efficient formulation is given.
+[2023b] that we used for our implementation; and we validated the model using the data from Yang et al.
+[2019].
+The proposed sigmoid-based TTC model, while simplified compared to multi-cue approaches Tian et al.
+[2023, 2025a], captures essential gap acceptance behaviors observed in empirical studies.
+Research on pedestrian crossing decisions reveals that time-to-arrival estimates—closely related to TTC—serve as primary decision variables, though these estimates are systematically distorted by approach speed and individual differences Raghuram Kadali et al.
+[2014].
+Specifically, pedestrians provide higher TTC estimates for higher approach speeds and accept smaller gaps accordingly, with age-related variations in conservativeness.
+Our sigmoid function’s adjustable parameter $c$ accommodates these individual differences, allowing the model to represent cautious to aggressive crossing behaviors within a unified framework.
 For the explicit motion model, it is assumed that the pedestrian’s choice of speed at the next time step can be modeled by
-$\dot{y}_{\mathrm{ped}}(t)=\frac{1}{1+\exp\left(-TTC(t)+c\right)}\cdot v_{\mathrm{ped}}^{\mathrm{ref}},$ (1)
-where c is a parameter that can be used to adjust the pedestrian’s characteristics, such as cautious or aggressive behavior. The $TTC$ is the time to collision and is computed by
-$TTC(t)=\frac{x_{\mathrm{ped}}(t)-x_{\mathrm{veh}}(t)}{v_{\mathrm{veh}}(t)}-\frac{y_{\mathrm{veh}}(t)-y_{\mathrm{ped}}(t)}{v_{\mathrm{ped}}^{\mathrm{ref}}}.$ (2)
+$\dot{y}_{\mathrm{ped}}(t)=\frac{1}{1+\exp\left(-TTC_{\mathrm{MPC}}(t)+c\right)}\cdot v_{\mathrm{ped}}^{\mathrm{ref}}.$ (1)
+The impact of the parameter $c$ is illustrated in Figure 2. The $TTC$ is computed as
+$TTC_{\mathrm{MPC}}(t)=\frac{x_{\mathrm{ped}}(t)-x_{\mathrm{veh}}(t)}{v_{\mathrm{veh}}(t)}-\frac{y_{\mathrm{veh}}(t)-y_{\mathrm{ped}}(t)}{v_{\mathrm{ped}}^{\mathrm{ref}}}.$ (2)
 The output function of (1) is a general sigmoid function and ranges between 0 and 1, which can be treated as the probability of the pedestrian crossing.
@@
 This model allows for real-time prediction of pedestrian motion and model-based interaction between the pedestrian and automated vehicle, making it simultaneously suitable for practical applications.
-The validity of (1) can be supported by the research of Kadali and Vedagiri (2020) and Forde and Daniel (2021), which propose also velocity-based models for pedestrian-vehicle interactions.
-The main benefit of the proposed model is that the decision layer is integrated into to control layer, thus the MPC can solve the trajectory planning and decision making tasks in one step.
+The model’s validity extends beyond its alignment with gap acceptance literature.
+By focusing on TTC as the core interaction metric, we follow established approaches in interaction-aware planning Crosato et al.
+[2024], while our probabilistic formulation provides advantages over deterministic models.
+The sigmoid mapping between TTC and crossing probability naturally handles uncertainty in pedestrian intentions—a critical aspect often overlooked in rule-based systems.
+Furthermore, validation against real-world trajectory data from Yang et al.
+[2019] demonstrates that this simplified model adequately captures crossing patterns in urban intersections, despite not explicitly modeling all perceptual cues that pedestrians might use.
+Importantly, the model’s simplicity enables real-time computation within MPC frameworks while maintaining behavioral plausibility.
+This trade-off between completeness and computational efficiency reflects broader challenges in social compliance for AVs, where perfect behavioral modeling must be balanced against practical deployment constraints Tian et al.
+[2025b].
+The main benefit of the proposed model (1) is that the decision layer is integrated into to control layer, thus the MPC can solve the trajectory planning and decision making tasks in one step.
+This can lead to a significant advantage compared to the state-of-the-art methods, where the decision-making and control layers are separated, which can lead to suboptimal solutions.
+Figure 2: The output of (1) with $v_{\mathrm{ped}}^{\mathrm{ref}}=1.4\frac{m}{a}$.
+Each curve represents a different pedestrian behavioral profile, with higher values of $c$ indicating more cautious decision-making and lower values reflecting assertive crossing behavior.
 Gap distance model (see e.g.
-(Tian et al., 2024)) inherently assumes that the pedestrian’s decision-making is based on the current gap value and assessing the safety without considering how that gap might change in the future.
+Tian et al.
+[2024]) inherently assumes that the pedestrian’s decision-making is based on the current gap value and assessing the safety without considering how that gap might change in the future.
 On the other hand, (1) modulates the interaction dynamics between pedestrian and vehicle, which evolves over time.
-It can account for variables such as speed, acceleration, and changing distances in a continuous manner and used for prediction interaction dynamics.
-Assuming a linear dynamics of the vehicle [^3] This assumption is commonly utilized by the application of autonomous vehicles, see e.g.
-Lynch and Park (2017, Chapter 13).
-and that the pedestrian moves in the $y$ direction, the following discrete dynamic system is obtained:
+It can account for variables such as speed, acceleration, and changing distances in a continuous manner and used for predicting interaction dynamics.
+Figure 3: Bird’s eye view of the scenario with the relevant distances for decision-making
+Assuming a linear dynamics of the vehicle - which is common by the applications of autonomous vehicles - and that the pedestrian moves in the $y$ direction, the following discrete dynamic system is obtained.
+The state vector is defined as $\boldsymbol{x}(t)=[x_{\mathrm{veh}}(t),\dot{x}_{\mathrm{veh}}(t),y_{\mathrm{ped}}(t),\dot{y}_{\mathrm{ped}}(t)]^{T}$, where $x_{\mathrm{veh}}$ and $\dot{x}_{\mathrm{veh}}$ are the position and velocity of the vehicle in the driving direction, respectively.
+The position and velocity of the pedestrian in the $y$ direction are denoted by $y_{\mathrm{ped}}$ and $\dot{y}_{\mathrm{ped}}$, respectively.
+The discrete time step is $\Delta t$.
+The explicit dynamics of the system is given by
 $\underbrace{\begin{bmatrix}x_{\mathrm{veh}}(t+\Delta t)\\ \dot{x}_{\mathrm{veh}}(t+\Delta t)\\ y_{\mathrm{ped}}(t+\Delta t)\\ \dot{y}_{\mathrm{ped}}(t+\Delta t)\end{bmatrix}}_{\boldsymbol{x}(t+\Delta t)}=\underbrace{\begin{bmatrix}1&{\Delta}t&0&0\\ 0&1&0&0\\ 0&0&1&\Delta t\\ 0&0&0&0\end{bmatrix}}_{{\mathbf{A}}}\underbrace{\begin{bmatrix}x_{\mathrm{veh}}(t)\\ \dot{x}_{\mathrm{veh}}(t)\\ y_{\mathrm{ped}}(t)\\ \dot{y}_{\mathrm{ped}}(t)\end{bmatrix}}_{\boldsymbol{x}(t)}$
-$+\underbrace{\begin{bmatrix}0.5\cdot{\Delta}t^{2}\\ \Delta t\\ 0\\ 0\end{bmatrix}}_{{\mathbf{B}}}u_{\mathrm{veh}}(t)+\underbrace{\begin{bmatrix}0\\ 0\\ 0\\ \frac{v_{\mathrm{ped}}^{\mathrm{ref}}}{1+e^{-TTC(t)+c}}\end{bmatrix}}_{\boldsymbol{z}(t)}$ (3)
+$+\underbrace{\begin{bmatrix}0.5\cdot{\Delta}t^{2}\\ \Delta t\\ 0\\ 0\end{bmatrix}}_{{\mathbf{B}}}u_{\mathrm{veh}}(t)+\underbrace{\begin{bmatrix}0\\ 0\\ 0\\ \frac{v_{\mathrm{ped}}^{\mathrm{ref}}}{1+e^{-TTC_{\mathrm{MPC}}(t)+c}}\end{bmatrix}}_{\boldsymbol{z}(t)}$ (3)
 where the desired acceleration of the vehicle $u_{\mathrm{veh}}(t)=a_{\mathrm{des}}$} is the system input.
-With the model (3.1), the prediction of the future states for $N$ steps is possible using the batch formulation Borrelli et al.
-(2017):
-$\mathcal{A}=\begin{bmatrix}\mathbf{A}\\ \mathbf{A}^{2}\\ \vdots\\ \mathbf{A}^{N}\end{bmatrix}\hskip 8.53581pt\mathrm{and}\hskip 8.53581pt\mathcal{B}=\begin{bmatrix}\mathbf{B}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{AB}&\mathbf{B}&\ldots&\mathbf{0}\\ \vdots&\ddots&\ddots&\vdots\\ \mathbf{A}^{N-1}\mathbf{B}&\ldots&\mathbf{AB}&\mathbf{B}\end{bmatrix}\hskip 5.69054pt\mathrm{and}\hskip 8.53581pt\mathcal{Z}=\begin{bmatrix}\mathbf{1}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{A}&\mathbf{1}&\ldots&\mathbf{0}\\ \vdots&\ddots&\ddots&\vdots\\ \mathbf{A}^{N-1}&\ldots&\mathbf{A}&\mathbf{1}\end{bmatrix}.$
+With the model (3.1), the prediction of the future states for $N$ steps is possible using the batch formulation, see Borrelli et al.
+[2017].
+The matrices $\mathcal{A}$, $\mathcal{B}$ and $\mathcal{Z}$ are defined as:
+$\mathcal{A}=\begin{bmatrix}\mathbf{A}\\ \mathbf{A}^{2}\\ \vdots\\ \mathbf{A}^{N}\end{bmatrix}\hskip 8.53581pt\mathrm{and}\hskip 8.53581pt\mathcal{B}=\begin{bmatrix}\mathbf{B}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{AB}&\mathbf{B}&\ldots&\mathbf{0}\\ \vdots&\ddots&\ddots&\vdots\\ \mathbf{A}^{N-1}\mathbf{B}&\ldots&\mathbf{AB}&\mathbf{B}\end{bmatrix}\hskip 5.69054pt\mathrm{and}$
+$\hskip 8.53581pt\mathcal{Z}=\begin{bmatrix}\mathbf{1}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{A}&\mathbf{1}&\ldots&\mathbf{0}\\ \vdots&\ddots&\ddots&\vdots\\ \mathbf{A}^{N-1}&\ldots&\mathbf{A}&\mathbf{1}\end{bmatrix}$.
 The prediction of the future state and input vectors are
@@
 These quadratic cost functions are
-Figure 2: Bird’s eye view of the scenario with the relevant distances for decision-making
 $J_{\mathrm{com}}$ $=\mathrm{w}_{\mathrm{com}}\cdot\boldsymbol{u}^{2}_{s}$ (6a)
@@
 $a_{\mathrm{veh}}^{\mathrm{min}}\leq u(t)\leq a_{\mathrm{veh}}^{\mathrm{max}},$ (7e)
-which computed the target acceleration of the AV.
-The optimization (7) is solved by a Python API of CasADi (Andersson et al., 2019), where the nonlinear programming solver with an interior point optimizer is applied, which can handle both the state (7c), (7d) and input constraints (7e).
+which computes the target acceleration of the AV.
+The optimization (7) is solved by a Python API of CasADi Andersson et al.
+[2019], where the nonlinear programming solver with an interior point optimizer is applied, which can handle both the state (7c), (7d) and input constraints (7e).
 With these hard constraints, the safety of the IAMPDM can be formally ensured by strictly enforcing operational boundaries.
@@
 Yang et al.
-(2022)) and should be taken into account to extend the MPC for the decision-making.
-Therefore, the crossing intention $I_{\mathrm{ped}}(t)$ of the pedestrian is introduced as a function of time and pedestrian behavior, which can vary between 00 and $1$, which can be used for predicting the actions of the pedestrian.
+[2022]) and should be taken into account to extend the MPC for the decision-making.
+Therefore, the crossing intention $I_{\mathrm{ped}}(t)$ of the pedestrian is introduced as a function of time and pedestrian behavior, which can vary between 0 and $1$, which can be used for predicting the actions of the pedestrian.
 Note that the aim of this paper is not the development of such intention-detection algorithms.
-It is assumed for this work that they are given since intention-detection algorithms can be found in the literature (Razali et al., 2021), (Rasouli et al., 2022), (Azarmi et al., 2023), (Zhou et al., 2024).
+It is assumed for this work that they are given since intention-detection algorithms can be found in the literature Razali et al.
+[2021], Rasouli et al.
+[2022], Azarmi et al.
+[2023], Zhou et al.
+[2024].
 To integrate the pedestrian’s crossing intention into the MPC formulation, the following two extensions are implemented:
@@
 As a solution, a discount function of the intention is proposed motivated by game theoretical applications, see e.g.
-Marín-Solano (2021):
+Marín-Solano [2021]:
 $I^{*}_{\mathrm{ped}}(t)=I_{\mathrm{ped}}(t_{0})\cdot 0.9^{K_{\mathrm{d}}\cdot t},$ (10)
@@
 If either the pedestrian or the vehicle has passed the intersection, no interaction occurs, allowing the vehicle to proceed at its reference speed, which is controlled by a simple velocity-tracking controller.
-1 Input: $ped$, $veh$, $I_{\mathrm{ped}}(t_{0})$
-2 Output: $veh\_acc$
+1Input: pedped, vehveh, $I_{\mathrm{ped}}(t_{0})$
+2 Output: veh_accveh\_acc
 3
-4if $is\_ped\_passed$ or $is\_veh\_passed$ then
+4if is_ped_passedis\_ped\_passed or is_veh_passedis\_veh\_passed then
 5 return velocity_control($ped,veh$)
 6else
-7       if $is\_discounting\_intention$ then
+7 if is_discounting_intentionis\_discounting\_intention then
 8      use (10)
@@
 13   Update vehicle state (4)
-14       Update $J_{\mathrm{MPC}}$ (6)
-15       $veh\_acc\leftarrow$ Solve (7)
-16       return $veh\_acc$
-17
-Algorithm 1 The IAMPDM Algorithm
-Note that the IAMPDM can be handle also.
-Managing dense pedestrian traffic is indeed a challenge in pedestrian crossing scenarios.
-In scenarios with many pedestrians crossing simultaneously, the IAMPDM framework can be adapted by interacting with the closest person.
-Another solution, is that the person with the highest estimated crossing intention is taken into account for the interaction.
-For both cases, the core working principle of the IAMPDM does not change.
+14 $veh\_acc\leftarrow$ Optimze (7)
+15 return veh_accveh\_acc
+16
+Algorithm 1 The IAMPDM Algorithm with MPC and Intention Integration
+Note that the main benefit of the IAMPDM is that its core structure remains the same across scenarios with few or many pedestrians.
+The only difference is the number of pedestrian models that need to be included in the MPC formulation.
+The complete IAMPDM algorithm is summarized in Algorithm 1.
 4 Experiment with the Interaction-Aware Model Predictive Decision-Making
@@
   • 2.
-    a rule-based decision-making (RBDM) introduced in Varga et al. (2023a).
+    a rule-based decision-making (RBDM) introduced inVarga et al. [2023a].
 The independent variable of the study design was the selection of the negotiation algorithm (MPDM, RBDM, NIA).
-Note: The Ethical Commission of WIVW approved the study based on Code of Ethics at WIVW – Summary for project proposals and articles/papers.
+For the experiment, N = 24 participants, (12 female and 12 male) were recruited from the WIVW Panel.
+Their mean age was 47.48 years (SD = 16.87).
+All participants were over 18 years old and had a valid driver’s license.
+A compensation of 25 Euros was paid for taking part in the experiment.
+The Ethical Commission of WIVW approved the study based on Code of Ethics at WIVW – Summary for project proposals and articles/papers.
+4.1 Description of the Compared Decision-Making Algorithms
+4.1.1 Non-Interactive Algorithm (NIA)
+The NIA is a conservative, non-interactive baseline that ignores explicit pedestrian crossing intentions; it applies a simple rule to stop the vehicle whenever a pedestrian is detected in the near or crossing zones (see Figure 3) and the time-to-collision (TTC) falls below a predefined threshold.
+which can lead to unnecessary stops and reduced traffic-flow efficiency compared to intention-aware methods such as IAMPDM and RBDM.
+4.1.2 Rule-Based Decision-Making (RBDM)
+The RBDM is a decision-making algorithm that considers the pedestrian’s intention to cross the street.
+The RBDM is based on a set of rules that determine the vehicle’s behavior based on the pedestrian’s position, speed, and intention to cross the street.
+The RBDM is designed to stop the vehicle if the pedestrian is in the safe zone and the TTC is less than a predefined threshold, but it also considers the pedestrian’s intention to cross the street.
+If the pedestrian is in the near zone and has a high intention to cross, the RBDM will slow down the vehicle to allow the pedestrian to cross safely.
+The RBDM is implemented as a comparison algorithm for the IAMPDM.
+For more details, see Varga et al.
+[2023a].
 4.2 Experimental Simulator Platform
@@
 In contrast, using a VR headset could cause simulator sickness, and the lack of visual representations of body parts in the virtual environment might lead to unnatural movement behavior.
-Therefore, we decided against a VR experiment based on these considerations and the advice of the ethical commission.
-of WIVW.
+Therefore, we decided against a VR experiment based on these considerations and the advice of the ethical commission of WIVW.
 The experimental setup includes
   • 1
-    A simulation computer with SILAB® [^5] SILAB is a software product of WIVW - Würzburg Institute for Traffic Sciences GmbH (www.wivw.de). software framework,
+    A simulation computer running SILAB® [^2] SILAB is a software product of WIVW - Würzburg Institute for Traffic Sciences GmbH (www.wivw.de) software framework that enables real-time traffic simulation with customizable vehicle behaviors and environmental conditions.
+The software provides accurate physics-based vehicle dynamics and allows integration of different decision-making algorithms for autonomous vehicles.
   • 2
-    A short-distance projector to display the scenario with the automated vehicle,
+    A short-distance projector mounted on the ceiling to display the scenario with the automated vehicle.
+The projector provides a large field of view (approximately 120 degrees horizontal) at a resolution of 1920x1080 pixels, enabling realistic perception of vehicle approach speeds and distances.
+The projection is calibrated to match real-world scale ratios.
   • 3
-    A body sensor to track the position and velocity of the test subjects and
+    A body sensor system consisting of HTC Vive Trackers to track the position and velocity of the test subjects in real-time with sub-centimeter accuracy.
+The tracking system operates at 90 Hz and provides position data of the test subjects that is synchronized with the simulation environment.
   • 4
     An input joystick for detecting the crossing intention of the test subjects.
-The advantage of using a joystick as input for the crossing intention helps to overcome the errors of computer vision algorithms that would be necessary to recognize the crossing intention of the pedestrian.
+The joystick provides an input signal ranging from 0 to 1 that represents the strength of the crossing intention.
+This direct input method avoids the complexity and potential errors of computer vision-based intention recognition systems while still enabling natural interaction dynamics.
+Using a joystick to signal crossing intention avoids computer‑vision errors while preserving natural interaction.
 Since this work does not attempt to validate such detection algorithms, the use of the joystick is a reasonable solution.
 Therefore, the experiment is even more rigorous in its objective of comparing different decision-making algorithms in a closed-loop manner.
-Figure 3: The simulator setup. Note that the light was turned off during the experiments, and the room was completely dark.
+Figure 4: The simulator setup with the marked street, where the pedestrian crosess.
+Note that the light was turned off during the experiments, and the room was completely dark.
+The only light source was the projector.
+The test subject wore the Vive Tracker on the waist to track their position and velocity.
+The joystick in the test subject’s hand was used to indicate their crossing intention.
 4.3 The Setups of the Decision-Making Algorithms
@@
 The inclusion of human preferences into IAMPDM tuning is one challenging aspect since it helps to enhance public acceptance of such technical systems.
-An automated tuning framework for pedestrian-AV interaction is presented in Varga et al. (2023a), which utilizes the optimization
+An automated tuning framework for pedestrian-AV interaction is presented in Varga et al. [2023a], which utilizes the optimization
 $\boldsymbol{\theta}^{*}$ $=\mathrm{arg}\,\mathrm{min}\,J_{\mathrm{glob}}$ (11a)
-[…]
+$\mathrm{s.t.}\;$ (7) (11b)
 to find the optimal parameter vector $\boldsymbol{\theta}$ of an intention-aware decision-making. The global cost function has the form
-$J_{\mathrm{glob}}=\int_{t_{0}}^{T_{\mathrm{end}}}k_{1}\cdot t+k_{2}\cdot\left|a^{2}_{\mathrm{max,veh}}(t)\right|-k_{3}\cdot\left|d_{\mathrm{min}}\right|+k_{4}\cdot\frac{1}{TTC(t)}\,\mathrm{d}t,$
+$J_{\mathrm{glob}}=\int_{t_{0}}^{T_{\mathrm{end}}}k_{1}\cdot t+k_{2}\cdot\left|a^{2}_{\mathrm{max,veh}}(t)\right|$
+$-k_{3}\cdot\left|d_{\mathrm{min}}\right|+k_{4}\cdot\frac{1}{TTC_{\mathrm{MPC}}(t)}\,\mathrm{d}t,$ (12)
 from which the optimal parameters of the intention-aware decision-making are obtained.
@@
 1
-2while Overaall result not satisfactory do
+2while Overall results not sufficient do
 3   Setting $k_{i},i=\{1,2,3,4\}$
@@
 As a result, distinguishing was simplified without using numbers, resulting in fewer matching errors by the test subjects.
-Since existing research does not address the human-in-the-decision-loop nature of pedestrian-AV interactions in unsignalised crossings, a key challenge was to instruct test subjects in a way that 1) facilitates interaction but 2) does not result in a fixed sequence of human actions that entirely predetermines their behavior.
+Since existing research does not address the human-in-the-decision-loop nature of pedestrian-AV interactions in unsignalized crossings, a key challenge was to instruct test subjects in a way that 1) facilitates interaction but 2) does not result in a fixed sequence of human actions that entirely predetermines their behavior.
 This balance is challenging, as too much freedom leads to high variability, whereas fully predefined behavior compromises the evaluation of the IAMPDM.
@@
 This challenge was addressed by instructing the test subjects to mimic altering their decisions: They started to cross the street, but they waited on the roadside.
-Therefore, we defined four scenarios, which describe approximately how the test subjects should behave, but no fixed sequence of human actions is defined, no, crossing time and distance specifications for specific actions are given in advance.
+Therefore, we defined four scenarios, which describe approximately how the test subjects should behave, but no fixed sequence of human actions is defined, no crossing time and distance specifications for specific actions are given in advance.
 The detailed instructions are given in the Appendix.
@@
   • H10
-    The use of intention-aware decision-making in autonomous vehicles significantly does not alter the crossing times and the criticality of intersection scenarios.
+    The use of intention‑aware decision‑making does not significantly alter the crossing times and the criticality of intersection scenarios.
   • H1w
@@
     The usage of intention-aware decision-making leads to a higher user preference in the intersection scenarios.
-To measure our hypotheses, we used three evaluation metrics defined in Westhofen et al. (2023):
+To measure our hypotheses, we used three evaluation metrics defined in Westhofen et al. [2023]:
   • 1.
@@
     $TTC(t)=\frac{y_{\mathrm{ped}}(t)+x_{\mathrm{veh}}(t)}{\mathrm{max}(v_{\mathrm{veh}}(t),\kappa)},$ (13)
-    where $\kappa=0.05$ in order to ensure the numerical stability, in cases of the pedestrian or the vehicle ar at a standstill. From that the average $TTC_{\mathrm{avg}}$ is calculated by
+    where $\kappa=0.05$ in order to ensure the numerical stability, in cases of the pedestrian or the vehicle are at a standstill. From that the average $TTC_{\mathrm{avg}}$ is calculated by
     $TTC_{\mathrm{avg}}=\frac{1}{N}\int_{t_{0}}^{T_{\mathrm{end}}}TTC(t)\,\text{d}t.$ (14)
@@
 Table 2: Answer Table for the test-subjects
-Note hypothesis H1 is evaluated by the objective metrics $TTC$, $DST$ and $T_{\mathrm{end}}$. For the assessment of hypothesis H2 the final questions are used.
-5 Results and Discussion
+Note hypothesis H1 is evaluated by the objective metrics $TTC$, DSTDST and $T_{\mathrm{end}}$. For the assessment of hypothesis H2 the final questions are used.
+5 Results
 This section presents the results of our experiment and provides an in-depth discussion.
@@
 Dash et al.
-(2023).
+[2023].
 5.1 Objective Results
@@
 First, the Kruskal-Wallis test is utilized for evaluating H1, because all three samples are compared, see e.g.
-Dalgaard (2008).
+Dalgaard [2008].
 The degrees of freedom of this test are $df=2$, and the significance level is chosen to $\alpha=0.01.$ Its null hypothesis is that there is no difference between the three decision-making algorithms.
@@
 In addition, Mann–Whitney–Wilcoxon tests are carried out to compare IAMPDM and RBDM.
-As it can be seen from Table 5, aside from $TCC$ in Delayed Remaining case, there is no statistical difference between IAMPDM and RBDM.
+As it can be seen from Table 5, aside from $TTC$ in Delayed Remaining case, there is no statistical difference between IAMPDM and RBDM.
 (a) Average crossing times
@@
 (b) Average DSTs
-(c) Average TTCs for Delayed Crossing
+(c) Average TTCs for Delayed Remaining
 Figure 6: Results in case of Delayed Remaining
@@
 $p_{\mathrm{Rem_{DST}}}$ $0.08$
-[…]
+$p_{\mathrm{Rem_{TTC}}}$ $9.47\text{\times}{10}^{-6}$
 Table 5: The resulting $p$-values of Mann–Whitney–Wilcoxon tests to compare IAMPDM and RBDM
@@
 The Kruskal-Wallis test is conducted to test the difference for statistical significance.
-The result is
-$\mathcal{H}_{\mathrm{Subj}}=14.56,$
-which indicates that the choice between intention-aware and non-interactive decision-making algorithms leads to statistically significant differences in the preference of the human test subjects.
-For a comparison between the IAMPDM and RBDM, a Mann–Whitney–Wilcoxon test is carried out, which yields
-$p_{\mathrm{Subj}}=0.154,$
-indicating that the difference between IAMPDM and RBDM is not significant.
+The result is $\mathcal{H}_{\mathrm{Subj}}=14.56,$ which indicates that the choice between intention-aware and non-interactive decision-making algorithms leads to statistically significant differences in the preference of the human test subjects.
+For a comparison between the IAMPDM and RBDM, a Mann–Whitney–Wilcoxon test is carried out, which yields $p_{\mathrm{Subj}}=0.154,$ indicating that the difference between IAMPDM and RBDM is not significant.
 Preferences NIA 5 RBDM 14 IAMPDM 5 Table 6: Preferences of the test subjects based on Q1 Final Question Score $\mu_{\mathrm{NIA}}$ 7.04 $\sigma_{\mathrm{NIA}}$ 3.629 $\mu_{\mathrm{RBDM}}$ 11.04 $\sigma_{\mathrm{RBDM}}$ 3.665 $\mu_{\mathrm{IAMPDM}}$ 10.00 $\sigma_{\mathrm{IAMPDM}}$ 2.690 Table 7: Mean values and standard deviations of Q2
@@
 5.3 Discussion and Limitations
-One of the limitations of our experimental setup lies in the assumption of perfect crossing intention detection via joystick input.
-While this simplified approach enabled controlled testing, real-world intention recognition involves inherent uncertainties due to sensor noise, occlusions, and ambiguous pedestrian behavior.
-It is necessary to highlight that higher TTC values are associated with more natural driving behavior of the AV.
+After presenting the results, we discuss the implications and limitations of this study.
+Firstly, it is necessary to highlight that higher TTC values are associated with natural driving behavior of the AV.
 This preference is reflected in the test subjects choosing the RBDM and IAMPDM over the more conservative NIA algorithm.
@@
 A closer examination of the objective results reveals that the intention-aware decision-making algorithms consistently outperform the NIA in all aspects.
-However, there is no significant difference between IAMPDM and RB, except in the average TTCs concerning Delayed Remaining (see Table 5).
+However, there is no significant difference between IAMPDM and RBDM, except in the average TTCs concerning Delayed Remaining (see Table 5), where IAMPDM exhibited slightly superior performance.
+For the other cases and measures, no significant differences between IAMPDM and RBDM were observed in this study.
 Additionally, we observed sporadic instances of extremely high crossing times for IAMPDM in Figure 6(a).
 These outliers primarily occurred when pedestrians hesitated mid-crossing, triggering repeated velocity adjustments of the AV.
-This reactive behavior suggests the need for better anticipation of pedestrian motion variability during the prediction.
-For the other cases and measures, no significant differences between IAMPDM and RB were observed in this study.
-These results suggest that a simple algorithm can achieve outcomes comparable to more complex methods in simple urban traffic scenarios.
-This implies that distinguishing between simple scenarios and complex scenarios is necessary for the development and validation of interaction-aware algorithms for autonomous vehicles.
-A simple scenario is defined as dyadic interactions with single pedestrians in controlled crossing situations, characterized by predictable trajectories, as in our experiment setup.
-Complex scenarios involve multiple interacting agents, mixed traffic modes, and ambiguous navigation goals.
-This distinction is crucial because while rule-based systems suffice for simple cases, complex scenarios demand the adaptive reasoning capabilities of model-based approaches like IAMPDM.
-The model (1) is also designed for more complex scenarios.
-It can extend to situations with multiple interacting agents.
-This flexibility can make it more applicable to a broader range of pedestrian-vehicle interactions, including those involving dynamic or unpredictable behaviors.
+This reactive behavior suggests the need for better anticipation of pedestrian motion variability during the prediction, see e.g.
+Tian et al.
+[2025a].
+Road traffic interactions involve reciprocal behaviors where agents adapt based on mutual observations.
+While Crosato et al.
+[2023] and Markkula et al.
+[2020b] use dynamical models that integrate social awareness for decision -making, our work aligns with their frameworks by modeling pedestrian intentions within an MPC structure.
+While our work utilizes TTC as a core interaction metric, it is recognized that pedestrians rely on a multitude of cues, see e.
+g.
+Tian et al.
+[2023].
+Furthermore, models like the one proposed by Tian et al.
+[2025a] validate this multi-cue hypothesis.
+Our IAMPDM can be extended with these additional cues in future work.
+The results suggest that a simple algorithm, like the RBDM, can achieve outcomes comparable to more complex methods in simple urban traffic scenarios.
+This implies that distinguishing between simple scenarios and complex scenarios is necessary for the validation of interaction-aware algorithms for autonomous vehicles.
+This distinction is crucial because while rule-based systems like RBDM suffice for simple cases, complex scenarios demand the capabilities of model-based approaches like IAMPDM.
+RBDM is a heuristic algorithm that adjusts velocity based on TTC and intention thresholds (e.g., yielding if pedestrian intention exceeds 0.5), monitoring pedestrian position and velocity.
+In contrast, IAMPDM employs a predictive model to optimize vehicle trajectories, considering pedestrian intentions and future states.
+This allows IAMPDM to anticipate pedestrian actions and plan accordingly, which is particularly beneficial in complex scenarios with multiple agents or ambiguous situations.
 Moreover, the subjective results offer interesting insights from the human-in-the-decision-loop experiment.
-The test subjects, who preferred the NIA, indicated that a more conservative approach could increase safety and be more favored by pedestrians.
-They had the comment such
+The test subjects who preferred the NIA indicated that a more conservative approach could increase safety and be more favored by pedestrians.
+They had comments such as:
   • 1.
@@
     Decision-making for the pedestrians pleasant.
-Most of the test subjects liked the RBDM more and criticized the long waiting time of the NIA.
-An interesting result is that the RBDM leads to better subjective results compared to IAMPDM: Most of the test subjects preferred RBDM over the IAMPDM, see Table 6.
+Most of the test subjects preferred RBDM over the IAMPDM, see Table 6.
 However, there is no significant difference in the objective results; see Table 7.
-Furthermore, most of them were not able not distinguish between RBDM and IAMPDM.
-They left comments like
+Furthermore, most of them were not able to distinguish between RBDM and IAMPDM, which are reflected in comments such as
   • 1.
@@
     [In case of RBDM:] no big difference to [IAMPDM]
-The reason for this result is the tuning framework, since the human factors component was tuned for both RBDM and IAMPDM in a similar manner, consequently, the subjective results were also similar.
+The reason for not detecting a significant difference between RBDM and IAMPDM is assumed to be the tuning framework.
+Both IAMPDM and RBDM were tuned with the same semi-automated, expert-in-the-loop process, which may reduce separability between methods and introduce bias toward smoothness and comfort.
+The NIA baseline relies on a waiting-time parameter that influences both safety margins and delay.
 Some general comments were that the test subjects were not sure whether the vehicle recognized them or not.
-To overcome this, implementing a human-machine interface (HMI) for a vehicle to communicate with pedestrians could significantly enhance trust.
+To overcome this, implementing an eHMI for a vehicle to communicate with pedestrians could significantly enhance trust.
 When a pedestrian sees a clear signal from the vehicle indicating it has detected them, it reduces uncertainty and potential accidents.
-The findings suggest that simply increasing the complexity of intention-aware decision-making algorithms does not necessarily lead to better performance or greater pedestrian preference.
-Additionally, the conditions under which these more complex and reliable model-based decision-making algorithms can achieve a safe scenario need to be validated in further studies.
+Results indicate that increased algorithmic complexity does not improve performance or user preference in simple, constrained scenarios.
+Preassumbly, advanced models like IAMPDM require targeted validation in more complex environments with larger interaction spaces to demonstrate their benefits.
+Future work will include additional metrics and testing in multi-agent, ambiguous settings to further explore these dynamics.
+The resulting practical implications for policy and deployment are that policymakers should set naturalness-aligned safety floors (e.g., min-TTC gaps and deceleration bounds), require conservative fallback when intention certainty is low, and mandate dual-objective calibration with fail-soft timeouts to prevent indecision oscillations.
 6 Summary
@@
 These findings highlight the significance of intention-aware decision-making for autonomous vehicles in urban settings.
-Our future work will feature an adaptive decision-making algorithm based on Varga (2024).
+Our future work will feature an adaptive decision-making algorithm based on Varga [2024].
 Furthermore, we plan to integrate a probabilistic estimation of human crossing intentions by modeling their stochastic behavior, see Škugor et al.
-(2023).
+[2023].
 CRediT authorship contribution statement
@@
 Ehsan Hashemi: Writing – review & editing, Investigation, Conceptualization.
-Acknowledgments
+Funding sources
 This work was supported by the Federal Ministry for Economic Affairs and Climate Action, in the New Vehicle and System Technologies research initiative with Project number 19A21008D.
@@
 The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
+During the preparation of this work, the authors used ChatGPT (OpenAI) to assist with improving language, readability, and organization of parts of the manuscript.
+After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
 Appendix A Instruction Materials